### v1

# Privacy-Preserving Dataset Combination

Abstract.
Access to diverse, high-quality datasets is crucial for machine learning model performance, yet data sharing remains limited by privacy concerns and competitive interests, particularly in regulated domains like healthcare. This dynamic especially disadvantages smaller organizations that lack resources to purchase data or negotiate favorable sharing agreements. We present SecureKL, a privacy-preserving framework that enables organizations to identify beneficial data partnerships without exposing sensitive information. Building on recent advances in dataset combination methods, we develop a secure multiparty computation protocol that maintains strong privacy guarantees while achieving ¿90% correlation with plaintext evaluations. In experiments with real-world hospital data, SecureKL successfully identifies beneficial data partnerships that improve model performance for intensive care unit mortality prediction while preserving data privacy. Our framework provides a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages. These results demonstrate the potential for privacy-preserving data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data resources. Our code is publicly available at https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Empirical scaling laws have established clear relationships between model performance and three key factors: compute, data, and model size (Hestness2017DeepLS, ; Kaplan2020ScalingLF, ). These relationships have driven remarkable improvements across computer vision, language processing, speech recognition, reinforcement learning, and healthcare applications (sun2017revisiting, ; mahajan2018exploring, ; brown2020language, ; hoffmann2022training, ; park2019specaugment, ; zhang2020pushing, ; ashvin2020accelerating, ; reed2022generalist, ; sheller2020federated, ; mckinney2020international, ). Beyond simple scaling, the diversity of training data has proven crucial for enhancing model robustness to distribution shifts and mitigating performance disparities across demographic groups (shen2024data, ; miller2021accuracy, ).
However, access to data varies significantly across entities and domains. While large tech companies have the data and compute resources to train the foundational models that now dominate general tasks, smaller players often lack such access. Domain-specific data is also becoming increasingly valuable for fine-tuning these general models (alsentzer2019publicly, ; gururangan2020don, ; lee2020biobert, ), creating a competitive advantage for the data owners. As a result, entities with domain-specific data are more reluctant to share it for free, opting instead to sell it in emerging data markets (acemoglu2022too, ; huang2021toward, ; liang2018survey, ). This dynamic particularly disadvantages smaller organizations, which often lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
Scaling datasets in regulated domains is particularly challenging due to legal constraints and unpredictable outcomes from altering the training data composition. In healthcare, for example, patient data is heavily regulated by laws such as the Health Insurance Portability and Accountability Act (HIPAA), which imposes strict data-sharing constraints to protect patient privacy. Moreover, accumulating data from multiple sources introduces the risk of domain shift, where data from different distributions may degrade model performance instead of improving it (koh2021wilds, ; wang2022generalizing, ; taori2020measuring, ). Contrary to the intuition that more data always leads to better performance, combining datasets does not guarantee improvements — in fact, performance can decrease when incorporating additional data sources (bradley2021unrepresentative, ; meng2018statistical, ; shen2024data, ). This non-monotonic behavior means that carefully selecting which datasets to combine is crucial, as using all available data may actually perform worse than using an optimal subset.
While existing methods aim to identify datasets combinations that can improve performance, they often assume access to all relevant datasets (shen2024data, ; ilyas2022datamodels, ). However, this is impractical because entities are reluctant to share their data due to privacy risks or the competitive value they associate with it. This reluctance to share creates a bottleneck in improving model performance. In high-stakes settings like healthcare, this bottleneck is particularly detrimental, as access to diverse data has the potential to drive significant advancements in patient care and outcomes (beam2018big, ; ghassemi2019practical, ; chen2020treating, ; beaulieu2019trends, ).
To address these challenges, we propose SecureKL, a privacy-preserving approach for guiding dataset combinations without requiring direct data or model sharing. Our work makes three primary technical contributions:
  1. (1)
    We introduce a framework for practical secure data combination that enables organizations to evaluate potential partnerships while maintaining data privacy. Our framework categorizes existing approaches by their privacy leakage risks and provides a systematic way to assess the trade-offs between data utility and privacy preservation.
  2. (2)
    We extend KL-XY score (shen2024data, ) with a secure multiparty protocol that enables privacy-preserving evaluation of potential data partnerships. Our protocol maintains strong privacy guarantees while utilizing the complete underlying datasets, achieving ¿90% correlation with plaintext computations.
  3. (3)
    Through extensive evaluation in low-data, high-stakes settings, we demonstrate that our method successfully identifies beneficial data partnerships for intensive care unit (ICU) mortality prediction, improving classifier performance for the source hospital.
  4. (4)
    In simulated scenarios requiring selection of three partner hospitals, Private-KL-XY outperforms alternative selection strategies including demographic-based selection (using gender, race, and age), and limited-sample plaintext selection.
We argue that our method presents an appealing trade-off of privacy and utility by preserving privacy for both parties while using all the underlying data.
2. Problem Setup
Consider a binary prediction task for ICU patient mortality based on electronic medical records. A source hospital $H_{o}$ has historical patient data $\mathcal{D}_{o}$ containing static past patient characteristics, prior medical records, and ICU outcomes. Other hospitals $\{H_{i}\}$ each has their patient data: $\{\mathcal{D}_{i}\mid i\in[1..N]\}$.
For this binary prediction task, hospitals typically optimize for performance metrics, for example the area under the receiver-operating characteristic curve (AUC). Using only their data, $H_{o}$ can train a model $\mathcal{M}$ with parameters $\theta$ to achieve:
(Baseline Performance) $\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$
where $f$ is their chosen algorithm with parameter $\theta$.
When $H_{o}$ has exhausted their own internal data, they may benefit from incorporating additional target data sources $T\subset[1..N]$. By combining datasets, i.e., $\mathcal{D}_{T}=\{\mathcal{D}_{i}\mid i\in T\}\cup\mathcal{D}_{o}$, $H_{o}$ can potentially achieve better results:
(Combined Performance) $\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$
We define the potential improvement from data addition as $\delta_{T}=\delta_{(o,T)}=\mathrm{AUC}_{T}-\mathrm{AUC}_{o}$. To add a single additional data source by setting $T=\{i\}$, the improvement is $\delta_{i}=\delta_{(o,i)}=\mathrm{AUC}_{i}-\mathrm{AUC}_{o}$. This leads to our central question: Without seeing target data, how does a hospital ascertain potential data sources to combine with?
Formally, given $n\leq N$, we seek a strategy $\pi$ that selects $n$ target datasets $T=\pi(\mathcal{D}_{o},n)$ to maximize model utility:
(Ideal Dataset Combination) $\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset{[1...N]\choose n}}\mathrm{AUC}_{T}$
Practical Considerations.
Computing every subset $T\subset{[1...N]\choose n}$’s associated $\delta_{T}$ is exponential in $n$. To make this problem tractable, we make two key assumptions. First, we apply strategies greedily, selecting top-ranked target datasets. With the ultimate objective of improving the source hospital’s prediction task, we fix $H_{o}$; to compare the trade-offs between strategies in Section 3, we apply each $\pi$ greedily to select top-$n$ institution(s) for $H_{o}$ without replacement. Second, in in data constrained settings, we aim to maximize the probability of positive improvement: $P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)$.
Kullback–Leibler Divergence.
Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work (shen2024data, ). KL divergence (kullback1951information, ), also called information gain (quinlan1986induction, ), describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
(Kullback–Leibler Divergence) $\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$
Because computing KL-divergence on datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial,  (shen2024data, ) proposes two groups of scores to make this divergence approximation tractable from small samples. Specifically, score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates — with the goal of inferring dataset membership. Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
(KL-XY Score) $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$
Details are described in Section 3.
Privacy Model for $\pi_{p}$.
We operate under a semi-honest privacy model—also known as honest-but-curious or passive security—where parties follow protocols but may probe intermediate values. Parties are “curious”, meaning that they can probe into the intermediate values to avoid paying for the data. This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation. This privacy preservation model incentivizes collaboration, improving upon methods in  (shen2024data, ) .
MPC Preliminary
To secure this divergence computation cryptographically, Secure Multiparty Computation (MPC) (yao1982protocols, ; shamir1979share, ) protocols are leveraged. Specifically, in $\mathrm{SecureKL}$, each party encodes $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ to preserve privacy for both parties. This is implemented with the research framework CrypTen (knott2021crypten, ), specialized for MPC and machine learning. Our algorithmic and engineering details are in Sections  3 and  4, respectively. For related secure techniques, see Section 5.4.
Additional Assumptions
Generally, we consider high stakes domains where disparate data may have additive benefits to the existing data. In order to make privacy boundaries tractable, we make the following additional assumptions:
  1. (1)
    Existing knowledge is not private. The hospitals are aware of each other having such data to begin with. The hospitals may know of the available underlying dataset size and format, which is assumed to be uniform across the hospitals in the setup to simulate unit-cost. Hospitals frequently know of each other’s resources, and the available ICU units are contentious, not kept secret.
  2. (2)
    Uniformity of $|\mathcal{D}_{i}|$. Though each hospital gets to price their data and set their own budget, for generality, the uniformity assumption allows us to use the number of additional data sources $n$ as the main ”budget proxy” across different strategies.
  3. (3)
    Legal risks of sharing any data are omnipresent in high stakes domains. The risks with sharing sensitive data in $\pi_{d}$ and $\pi_{s}$ are not made explicit, but assumed to be ”medium” and ”medium-to-high” respectively. This abstraction side-steps legal discussion, which would go beyond the scope of our paper.
  4. (4)
    No malice is assumed on any of the parties involved, as each hospital wants to authentically sell their data and set up a potential collaboration. This assumption becomes stronger when the number of parties grows or when the setup changes to potentially more competitive industries with less trust. We note our limitations in Section 4.5.
3. Methods
Figure 1. Our Method $\mathrm{SecureKL}(\mathcal{D}_{o},\mathcal{D}_{i})$. Each side encrypts their data, and a model is trained on their joint data. Using private KL-based measures, their distance is computed, and the final result is revealed after decryption, which requires both parties.
[…]
Data is encrypted on each side.
We summarize our data acquisition strategies differentiated by leakage risks, which correlate with potential costs:
Category 1, medium-to-high leakage, sharing raw data. $\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered (lds, ). Though leakage can be controlled through $k$, the data is inherently sensitive. The underlying distance uses  (shen2024data, ) ’s KL-XY Score.
Category 2, medium leakage, sharing summary statistics. $\pi_{d}(n)$ uses demographic metadata to guide data selection. This is implemented through ratio distance between source and target distributions, which may be considered aggregates therefore potentially not sensitive, such as when the underlying aggregation function $\phi$ is differentially private.
Category 3, zero leakage, sharing no additional information besides what is assumed public. There are two methods: a. Blind selection baseline: $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. Prior works suggests that when $n=1$, randomly selecting a source in hospital ICU may be risky and inefficient. b. Our method $\pi_{p}(n)$ selects data sources based on privacy-preserving measure for data combination, specifically Private KL-XY.
3.1. Trivial Baseline: Blind Selection
Blind selection refers to the process when no information is provided. $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. This random strategy may evade selection biases and help gather diverse data. Yet, prior work  (shen2024data, ) suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
3.2. Sharing Summary Statistics
A relaxation to sharing no sensitive data is to share metadata. While demographic traits are often causal and available, their exact cause in relation to the task is not a priori established (without a highly effective model), therefore their success in distributional-matching is not guaranteed to be strong. Additionally, in practice, the most effective model that results from data combination may or may not be causally-sound. Nevertheless, we posit alternative strategy $\pi_{d}(n)$ to find the demographically close candidates to guide data selection: Let $\phi:\mathcal{D}\to\mathbb{R}^{m}$ be an $m$-dimensional summary statistic of a demographic feature i.e. the racial distribution of patients. Then, we use the distributional distance between $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, characterized by their $L_{2}$-distance through $\phi$:
(Demographic-based Strategy) $\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$
3.3. KL-based Methods, in Plaintext
$\pi_{s}(n,k)$ assumes each of the candidate hospitals will share a set of raw data. In ICU data, simulate that a default of $1\%$ is shared, so $k=3000\times 1\%=30$, though we run experiments with $k\in\{3,30.300,3000\}$ (Section 4.3). Though leakage can be controlled through $k$, the data is inherently sensitive.
This is implemented with KL-based measures similar to (shen2024data, ) . To recap, $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a ”metric” that satisfies triangle inequality. Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be ”close” to the target distribution $P_{i}$, but not the other way around:
(Ideal Estimator) $\mathrm{KL}(P_{o}||P_{i})=\int_{x\in\mathcal{X}}\log\frac{P_{o}(\mathop{}\!\mathrm{d}x)}{P_{i}(\mathop{}\!\mathrm{d}x)}P_{i}(\mathop{}\!\mathrm{d}x)$
Because we only have access to finite data $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, approximations are needed. Typically, a learned model can capture distributional information, used to estimate continuous entropy. Thus the joint distribution of features and labels from both the source and target are included, with the goal of deriving an efficient estimator for $\mathrm{KL}(P_{o}||P_{i})$ that captures distributional shift from source to target.
Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model (Cox1958TheRA, ) on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covaraiates – with the goal of inferring dataset membership. A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful.  (shen2024data, ) established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task. We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
(KL-X) $\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$
(KL-XY) $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$
We focus on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, and reproduce (shen2024data, ) ’s results that it is predictive of downstream change in AUC.
Let the score function $g_{\mathrm{KL}}$ be the approximate of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$. The strategy selects the most likely hospital with the closest distance under the measure:
(KL-based Strategy, in plaintext) $\pi_{s}(n=1,k=K)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{KL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$
When only a subset is available, this function is adjusted by swapping $\mathcal{D}_{i}$ for $\mathcal{D}_{i}^{\prime}\subseteq\mathcal{D}_{i}$ where $|\mathcal{D}_{i}^{\prime}|=k$. We denote the full dataset size as $K=|\mathcal{D}_{i}|$.
3.4. SecureKL: Private KL-based Method
Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public). Specifically we leverage the MPC based framework provided by CrypTen (knott2021crypten, ), a library designed for privacy-preserving machine learning, to implement private $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. As illustrated in Figure 1, the logistic regression as well as the scoring need to be implemented in private. Our code is publicly available  [^1] https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
Denote the private encoding of $x$ as $[x]$.
(Secure KL-X) $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$
(Secure KL-XY) $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}([x,y])).$
Let the score function $g_{\mathrm{SKL}}$ be the secure approximation of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$. The strategy selects the most likely hospital with the closest distance under the measure:
(SecureKL Strategy, encrypted) $\pi_{p}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{SKL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$
As shown in Figure 1, any KL-based measure $g_{\mathrm{SKL}}$ can be adapted to our setup. We mainly use $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ as the underlying measure. Its performance is detailed in Section 4.4. Additionally, even though our implementation measures distance of data between one source and one target party, the setup readily extends to accommodating multiple parties. We note the engineering limitations in Section 4.5.3.
4. Experiments
4.1. Experimental Setup
We validate our method and demonstrate its applicability using the eICU Collaborative Research Dataset (Pollard2018TheEC, ), which contains over $200,000$ admissions from $208$ hospitals across the United States. Following the data cleaning and exclusion criteria outlined by (Water2023YetAI, ) and (shen2024data, ), we selected the $12$ hospitals with the highest number of patient visits (each with at least $2000$ patients) as our $\mathbf{H}$. Each strategy would compute with the same $K=3000$ records, as the total available data per hospital.
We simulate the problem setup for each hospital with the $24$-hour mortality prediction task. The strategy comparisons described in Section 3 are implemented using $1500$ samples and the AUC is evaluated on $400$ samples for all of our experiments unless otherwise noted. This follows training and evaluation protocols in Yet Another ICU Benchmark (Water2023YetAI, ). For the data combination experiments that compute AUC change $\delta_{i}$ or $\delta_{T}$, to match  (shen2024data, ) , we take $1500$ random samples from each selected dataset and combine it with $1500$ samples from $\mathcal{D}_{o}$ (fixed across all experiments).
Implementing $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be privacy-preserving requires training logistic regression model in private. This is used in the private setting to estimate $g_{\mathrm{SKL}}$$\mathrm{Score}([X])$ or $\mathrm{Score}([X,Y])$ – for each pair of hospitals. Our experiments train encrypted logistic regression in CrypTen (knott2021crypten, ) using the library’s SGD optimizer. To ensure a fair comparison between the scores obtained through plaintext and encrypted settings, we re-implement plaintext Score($X$) and Score($X,Y$) using logistic regression with SGD in PyTorch (paszke2019pytorch, ). This is because encrypted version of L-BFGS – the optimizer prior work (shen2024data, ) uses in plaintext-only with Scikit-Learn (scikit-learn, )– is not available in CrypTen, though it leads to better downstream performance. Hyper-parameter tuning for SGD in private and plain text are performed independently, with the details in Appendix B.
4.2. Experimental Questions
We ask three sets of questions:
  1. (1)
    Consistency: Does using multiparty implementation sacrifice original measure’s effectiveness? Practically, we evaluate this through the analysis of private and plaintext scores. For our selected metrics, we expect AUC change to be negatively with KL-based measures—meaning the closer the additional target dataset is to the source hospital, the more the AUC will improve compared to other potential target datasets. Section 4.3 tests the correlation of our private scores and plaintext scores with full access (setting $k=K$). In addition to computing Spearman’s rank correlation coefficient of the KL-scores, we probe the discrepancy of the downstream effect between $\{\delta_{i}|i=\pi_{s}(n=1,k=K)\}$ using plaintext $\mathrm{KL}_{\mathcal{X}},\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ with all the underlying data and $\{\delta_{i}|\pi=\pi_{p}\}$ using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ for each source hospital $H_{o}\in\mathbf{H}$.
  2. (2)
    Positivity: Does our method pick hospitals that reliably improve performance? If source dataset $D_{o}$ can only add data from $n$ more hospitals, does our measure lead to eventual AUC improvements? In Section 4.4, we test our framework on a multi-dataset combination experiment and find that it successful improves the source hospital’s downstream outcome. Specifically, when selecting a single additional data source ($n=1$), all but 2 hospitals improves, and when selecting top 2 or 3, all hospitals see a positive AUC change. This shows a consistent added benefit.
    Lastly, Section 4.4.1 compare with different strategies proposed in Section 3, and Section 4.4.2 analyze the benefits of using private dataset combination $\mathrm{SKL}$.
  3. (3)
    Error analysis: If our privacy-preserving method is not the dominant strategy against alternatives including limited data accessibility, why is that? Section 4.5 performs additional analyses on (a) hospitals with low $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ correlations, and (b) hospitals lagging AUC improvements using the random strategy $\pi_{0}$ or limited-sample strategy $\pi_{s}$ for selecting $n=3$ candidate hospitals.
Lastly, we note engineering hurdles to scale to the real-world in Section 4.5.3. We hereby detail our results.
4.3. Consistency Between Plaintext and Encrypted Computations
Our encrypted computations are programmed with CrypTen, when the plaintext counterpart is PyTorch-only. We show using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ lead to highly comparable behavior as $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}}$, respectively.
Spearman’s Rank Correlation Coefficient for Underlying Scores
For each source hospital $H_{o}$, use all full samples for $\mathcal{D}_{i}$. Between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ on $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ for all remaining hospitals $H_{i}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.908$ with a range of $[0.691,1.0]$, obtaining $p<0.02$ across all hospitals. Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$. After applying Hochberg false discovery rate correction (Benjamini1995ControllingTF, ), our p-values remain significant. This range is an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations. When we unify the SGD hyperparameters, we indeed get a tighter range. For all 12 hospitals, see appended Appendix C for details.
Downstream Model Improvements
We further simulate the effect by adding encryption through its impact on the downstream AUC. This examines whether there will be a shift in the full hospital ranking, if we switch from a plaintext setup to encrypted. For $H_{o}\sim\mathbf{H}$, we measure $\delta_{i}$ that results from adding $\mathcal{D}_{i}$ to $\mathcal{D}_{o}$ for all $i$. This correlates all target hospitals $\{H_{i}\}$ with their ground truths $\{\delta_{i}\}$ in the case of picking a single target hospital. We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.182$ and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.184$ ($99\%$ matching). Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of $-0.164$ with $\delta_{i}$. For all strategies’ correlations with ground truth at $n=1$, see Appendix A.
Figure 2. AUC change $\delta_{T}$, after including top-$3$ hospital per strategy, over all source hospitals. Our private strategy $\pi_{p}$ is compared with (a) demographic-based $\pi_{d}$ for gender, age, and race. (b) plain text limited-sample $\pi_{s}(k=K)$ for $k=300$ (10%) and $k=30$ (1%).
4.4. Positivity of SecureKL and Practical Implications
Overall Positivity
We evaluate the practical utility of SecureKL by applying it in a multi-source data combination experiment, where $n\in\{1,2,3\}$. For $n=1$, we find that $\pi_{p}$ improves AUC in 10 out of the 12 hospitals. When $n=2$ and $n=3$, we find that using $\pi_{p}$ consistently improves AUC for all hospitals. Overall, 34 out of the 36 dataset combinations we evaluate on have an AUC improvement $\delta_{T}>0$, suggesting that $\pi_{p}$ is a reasonable strategy for selecting hospital dataset combinations with a high expected return $\mathbb{E}[P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)]$ for the source hospital from using our strategy.
4.4.1. Comparing With Alternative Strategies
Other strategies – $\pi_{0}$, $\pi_{d}$, and $\pi_{s}$ – can also arrive at positive datasets. Comparing private method $\pi_{p}(n=3)$ to other strategies at $n=3$, we find the following results, illustrated in Figure 2:
  1. (1)
[…]
  2. (2)
    Demographic-based strategies underperform compared to $\pi_{p}$ on average. However, we observe that $\pi_{d}$-gender can be highly effective for a subset of hospitals, as it achieves the highest 75th percentile (Q3) of $0.033$ among all strategies. This indicates that for 25% of hospitals, $\delta_{T}\geq 0.033$. Despite this, $\pi_{d}$-gender has a lower median value of $0.012$ compared to $\pi_{p}$, exhibits a high standard deviation ($0.022$), and degrades the performance for certain hospitals. Similarly, $\pi_{d}$-age has a median of $0.014$, and $\pi_{d}$-race has a median of $0.008$, both lower than $\pi_{p}$’s median.
  3. (3)
    Plaintext small-sample strategies, $\pi_{s}$, outperform all demographic-based methods but slightly underperform relative to $\pi_{p}$. For instance, $\pi_{s}(k=300)$ has a median $\delta_{T}$ of $0.0178$, and although it achieves $\delta_{T}>0$ across all hospitals, it performs worse on average compared to $\pi_{p}$ and exhibits a higher standard deviation ($0.017$). $\pi_{s}(k=30)$ has a median $\delta_{T}$ of $0.0165$. Compared to other strategies, it has the largest standard deviation ($0.024$), and it degrades the performance for some hospitals.
In summary, our method $\pi_{p}$ achieves the highest AUC improvement on average with the lowest standard deviation, demonstrating consistent improvement for all hospitals. In contrast, demographic-based and plaintext small-sample strategies exhibit greater variability, with some strategies improving performance for specific subsets of hospitals but underperforming or degrading results in others.
4.4.2. SecureKL Analysis
After establishing that $\pi_{p}$ with $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$is a robust strategy in practical downstream performance, we hereby synopsize the benefits of $\mathrm{SecureKL}$ and elaborate on their practical implications.
A Principled Approach To Data Minimization. Our major contribution is to match plaintext performance with no data sharing. Using MPC provides input privacy, meaning that if both hospitals only want to know the resulting score, the computation can be done without leaking original data. This strong guarantee can significantly ease the tension related to privacy and compliance in setting up a collaboration, leading to a practical ”data appraisal stage” in data-limited high stakes domains.
In the case where that output can be sensitive, i.e., when hospitals query each other multiple times and accrue information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response (dwork2014algorithmic, ).
In theory, any data combination method (if Turing-complete) can be made private; yet, in practice, balancing the right trade-off of utility and privacy is non-trivial. Barring engineering difficulties, not all algorithms readily adapt efficiently in private. Prior work  (xu2022data, ) included the trained model and test data in private; while relatively exact, complex methods would exacerbate the same operational limitations discussed in Section 4.5.3.
Gain from Data Availability. In contrast to limited-sample approaches, a key advantage for our method $\pi_{p}$ is that it takes advantage of all of the underlying data – generally impossible with non-secure methods for private data in heavily regulated domains. The general intuition is that data is localized; therefore, once a good target hospital is identified, we should acquire all of the data. It may be tempting to assert that we prefer the highest $k$ for data addition algorithms as well. In our experiments, while this is generally true, the smaller $k$ sometimes outperform larger $k$ in plaintext strategy $\pi_{s}$, which we investigate in Section 4.5 and in Figure 5. This occasionally non-monotonic behavior mirrors the challenge of data combination itself: even within one source dataset for the same estimator, more data is not necessarily better. This suggests a domain-specific alternative to sharing a large amount of data for some source hospital, and points to future directions to using secure computation on a minimal-sized sample dataset for minimal performance overhead while remaining private.
4.5. Error Analysis
4.5.1. Underlying Score Limitations
Data addition algorithms underpin the effectiveness of our method. Even if $\mathcal{D}_{o}$ obtains access to all the plaintext data, there is no guarantee that $\pi_{p}$ can correctly predict whether the data is useful. As seen in Figure 4, Hospital 243’s utility when acquiring another data set is badly correlated with plaintext and encrypted KL-XY scores. This leads to its bad strategy for acquiring the top 3 hospitals, as seen in the middle pane of Figure 5. Interestingly, for this hospital, no other informational strategy excels, either, so choosing a random 3 may be preferred.
This behavior stems from the underlying measure, not from adding secure computation: in Figure 4 and Figure 5, the encrypted performance closely follows that of plaintext performance, for both good and bad downstream correlations.
4.5.2. Sometimes, Not All Underlying Data Is Needed
Relatedly, when seeing a few samples can successfully identify useful candidate hospitals, $\pi_{p}$ does not always outperform $\pi_{s}$ on small samples.
In the right panel of Figure 5, hospital 199, the smaller sample sizes achieve a score that better reflects ground truth as a data addition strategy. In that case, the hospital may not need the full sample to know which target hospitals to collaborate with.
This behavior is specific to the interaction of the data and the underlying score, and does not affect the general insight that adding private computation preserves privacy (and eases privacy-related risks that hinder data sharing). We further note that our method still clearly applies to encrypted computation on a smaller data set under data minimization.
Figure 3. Hospital 243: Underlying KL-score performs poorly
[…]
NEED UPDATING some conclusion and description
Figure 4. Hospital 420: Underlying KL-score performs well
[…]
NEED UPDATING
Figure 5. Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{p}(k=30)$ and $\pi_{0}$. Middle: All strategies perform similarly. Right: $\pi_{p}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. In all panels, the bars represent the standard deviation.
[…]
NEED UPDATING
4.5.3. Encrypted Computation Limitations
Engineering a secure system for machine learning requires both machine learning and software engineering knowledge. We share our our code and method, but also note potential challenges with deploying our secure computation:
  1. (1)
    Operational: engineering personnel limitations. While our implementation requires little cryptographic knowledge to deploy, it still needs technically-trained staff at each participating hospital to collaborate and maintain.
  2. (2)
    Engineering: Extending any MPC protocol is non-trivial, as security engineering is a specialized skill. While $\mathrm{SecureKL}$ applies broadly to other underlying scores in multi-party setups, every new algorithm requires software engineering - prototyping, tuning, debugging —which can be especially costly for hospitals.
  3. (3)
    Framework Limitation: While CrypTen is designed to accommodate PyTorch, it is a research tool where not all plain text functionalities are implemented. Writing optimizers – such as L-BFGS – and custom operators that are not readily available requires both machine learning and cryptography knowledge.
  4. (4)
    Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly. As detailed in Appendix  B, our hyperparameters for SGD are indeed different in encrypted and plaintext settings. However, encrypted computation hides loss curves and training details by default, complicating development.
5. Related Work
5.1. Data Valuation and Pricing
The question of how to assess the impact and therefore worth of data has been well studied. Data valuation as a field seeks to quantify the contribution of individual data points or datasets to model performance. Shapley value-based approaches provide theoretically grounded valuations but scale poorly to large datasets (ghorbani2019data, ). More efficient methods include influence functions (koh2017understanding, ) and leave-one-out testing (cook1979influential, ). Recent work has extended these concepts to dataset-level valuation (jia2019towards, ). As practitioners create larger datasets, the emergence of data marketplaces has sparked interest in data pricing mechanisms (kumar2020marketplace, ). Query-based pricing (koutris2015query, ) and outcome-driven valuations (radic2024pricing, ) aim to balance seller compensation with buyer utility. While these approaches inform fair data exchange, they typically assume direct access to data, unlike our privacy-preserving method.
5.2. Alternative Approaches to Data Sharing
Recent work has explored several approaches to mitigate data sharing constraints while maintaining model performance. We discuss two primary directions: synthetic data generation and transfer learning from public pretraining.
Synthetic data generation
Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy. Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data (goncalves2020generation, ), medical imaging data (thambawita2022singan, ), and electronic health records (baowaly2019synthesizing, ). These methods can preserve statistical properties of the original data while providing differential privacy guarantees. Transforming data into a similar form that de-sensitizes certain attributes can be desirable (drechsler2011synthetic, ; howe2017synthetic, ; nikolenko2021synthetic, ; gonzales2023synthetic, ; sweeney2002k, ). Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself (jordon2021hide, ). Additionally, outside of the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks (narayanan2006break, ; jordon2021hide, ).
However, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables (goncalves2020generation, ). For tabular data, methods like CTGAN (xu2019modeling, ) and TVAE (xu2019modeling, ) have demonstrated ability to learn complex distributions while preserving correlations between features. However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance (assefa2020generating, ). Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees (jordon2018pate, ). While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset (yang2024tabular, ).
Public pre-training and private fine-tuning
Transfer learning via public pretraining has become increasingly popular for domains with limited private data access. BioBERT (lee2020biobert, ) and ClinicalBERT (alsentzer2019publicly, ) demonstrated that pretraining on PubMed abstracts and clinical notes can improve performance on downstream medical tasks. Similar approaches have emerged in other regulated domains, including FinBERT (liu2021finbert, ) for financial applications. However, the effectiveness of transfer learning depends heavily on domain alignment. One study showed that continued pretraining on domain-specific data significantly outperforms generic pretraining when domains differ substantially (gururangan2020don, ). This presents challenges for highly specialized fields where public data may not capture domain-specific patterns (le2022language, ). Recent work has explored methods to quantify and optimize domain adaptation. Adaptive pretraining strategies (howard2018universal, ) and domain-specific vocabulary augmentation (gururangan2020don, ) have shown promise in bridging domain gaps. However, these approaches still require substantial compute and may not fully capture specialized domain knowledge present in private datasets.
5.3. Benefits of Data Scaling Beyond Performance
Recent work has demonstrated that increasing dataset size and diversity yields benefits beyond raw performance metrics. Large-scale training data has been shown to improve model robustness to distribution shifts (taori2020measuring, ) and reduce demographic performance disparities (chen2018my, ). Studies of vision models trained on increasingly large datasets show improved out-of-distribution generalization (miller2021accuracy, ). Similarly, language models trained on diverse data demonstrate better performance across different domains and demographic groups (dodge2021documenting, ).
5.4. Secure and Confidential Computation
Secure and confidential computation encompasses cryptographic techniques that protect information privacy during computation. In a two-party setup between a model owner and data owner, these methods enable computing joint functions on private inputs without revealing them to other parties.
This requires an encoding scheme $\mathrm{Enc}(\cdot)$ that satisfies the homomorphic property: $\mathrm{Enc}(A)\circ\mathrm{Enc}(B)=\mathrm{Enc}(A\circ B)$, where $A$ and $B$ represent data held by two parties. The inverse function $\mathrm{Enc}^{-1}(\cdot)$ must exist to decode the final output: $\mathrm{Enc}^{-1}(A\circ B)=A\circ B$. Considering an “honest-but-curious” threat model, where parties aim to jointly compute on privately-held data, two main approaches emerge.
Fully Homomorphic Encryption (FHE)
FHE enables arbitrary additions and multiplications on encrypted inputs. While it represents the gold standard for encrypted computation, adapting it to modern machine learning is challenging due to computational constraints from growing ciphertext size. FHE implementations typically use lattice-based schemes requiring periodic “bootstrapping” (key refreshing and noise reduction through re-encryption) via methods like the CKKS scheme (cheon2017homomorphic, ). This introduces cryptographic parameters that non-experts struggle to configure effectively.
Secure Multi-party Computation (S-MPC)
SMPC enables multiple parties to compute functions over private inputs while revealing only the final output (yao1982protocols, ; shamir1979share, ). An MPC system uses key exchanges, encryption schemes, and secure communication to ensure only encrypted data leaves owner control (micali1987play, ). Though generally faster than FHE, SMPC’s engineering complexity and communication overhead can limit adoption. Traditional private training approaches that completely hide data can also impede essential model development tasks like inspection, monitoring, and debugging.
5.5. Secure Data Combination
Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance. Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data (aono2017privacy, ). However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets (mcmahan2017communication, ). More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment. These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets (leung2019towards, ; chakraborty2024privacy, ). Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information (duan2021flexible, ). Several systems have been developed to facilitate secure data combination in specific domains. In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance (raisaro2018m, ). Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality (liu2021finbert, ).
5.6. Other Privacy-Preserving Methods
Federated Learning.
Cross-silo federated, decentralized, and collaborative machine learning  (mcmahan2017communication, ; li2020federated, ; bonawitztowards, ; kairouz2021advances, ) focus on acquiring more data through improved data governance and efficient system design. Healthcare machine learning is considered especially suitable, as health records are often isolated (rieke2020future, ; xu2021federated, ; nguyen2022federated, ). Yet, even though no raw data is shared, model parameters or gradients flow through the system. As the federated computing paradigm offer no privacy guarantee, the system is vulnerable to model inversion (geiping2020inverting, ) and gradients leakage attacks (boenisch2023curious, ; zhu2019deep, ). A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimisation is traded off with privacy (lyu2022privacy, ; raynal2024conflict, ). Building on the insight that useful data is often disparately owned, we tackle the specific incentive problem between pairs of data players where one side trains the model, instead of scaling up a federation (number of parties) to address data access issues. We thus focusing on making this exchange efficient, accurate, and private.
Compared to vanilla Federated Learning, an MPC system (shamir1979share, ; yao1982protocols, ; bonawitz2017practical, ; knott2021crypten, ) provides stronger guarantee in terms of input security. Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements. Our work extends the line of works  (xu2022data, ; yang2024fedfed, ; bonawitz2017practical, ) that demonstrates the potential of incorporating MPC in various federated scenarios. On the practical side, unlike mobile-based networks for secure federated learning protocols (bonawitztowards, ), our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
Differential Privacy
Differential privacy (DP) (dwork2006differential, ) offers formal privacy guarantees for sharing data and training machine learning models. While DP mechanisms can protect individual privacy when releasing model outputs or aggregated statistics, they face significant limitations for inter-organizational data sharing. The primary challenge is that DP operates on already-pooled data, but organizations are often unwilling to share their raw data in the first place (dwork2014algorithmic, ). Even when organizations are willing to share data, the privacy guarantees of DP come at a substantial cost to utility, particularly in machine learning applications. DP-SGD, the standard approach for training deep neural networks with differential privacy, significantly degrades model performance compared to non-private training (abadi2016deep, ). This performance impact is especially pronounced in data-constrained settings, where recent work has shown that large models rely heavily on memorization of rare examples that DP mechanisms tend to obscure (feldman2020neural, ). The privacy-utility trade-off becomes even more challenging when dealing with high-dimensional data or complex learning tasks. Studies have demonstrated that achieving meaningful privacy guarantees while maintaining acceptable model performance requires prohibitively large datasets (bagdasaryan2020backdoor, ). This limitation is particularly problematic in specialized domains like healthcare, where data is inherently limited and performance requirements are stringent (geyer2017differentially, ). Recent work has attempted to improve the privacy-utility trade-off through advanced composition theorems and adaptive privacy budget allocation (papernot2021tempered, ). However, these approaches still struggle to match the performance of non-private training, especially when working with modern deep learning architectures (tramer2020differentially, ). While differential privacy offers important theoretical guarantees, our work focuses on the practical challenge of enabling data owners to evaluate potential partnerships before sharing any data, addressing a key barrier to collaboration that DP alone cannot solve.
6. Conclusion
Our work demonstrates that privacy-preserving data valuation can help organizations identify beneficial data partnerships while maintaining data sovereignty. Through SecureKL, we show that entities can make informed decisions about data sharing without compromising privacy or requiring complete dataset access. As the AI community continues to grapple with data access challenges, particularly in regulated domains like healthcare, methods that balance privacy and utility will become increasingly critical for responsible advancement of the field. As noted in Section 4.5, our approach has several limitations, including the fact that, despite impressive aggregate results, our method is less effective for individual hospitals; this finding is fertile ground for future work. Additionally, our work present opportunities for follow-up research. Our method assumes static datasets and may not generalize well to scenarios where data distributions evolve rapidly over time. A sequential version of our framework may more closely model dynamic data collaborations. Future work should explore extending these techniques to handle more complex data types and dynamic distribution shifts while maintaining strong privacy guarantees.
Appendix A Correlation with downstream performance
On Table 1, we report the Pearson correlations between $\pi_{s}(k=K)$ for $k\in\{3,30,300,3000\}$ and $\delta_{i}$. On Table 2, we report the Pearson correlations between different strategies and $\delta_{i}$.
k $\rho$ p-value
3 -0.063 4.70e-01
30 -0.082 3.47e-01
300 -0.059 5.00e-01
3000 -0.184 3.47e-02
k $\rho$ p-value
3 -0.158 7.02e-02
30 0.167 5.60e-02
300 -0.097 2.70e-01
3000 -0.284 9.47e-04
Table 1. $\rho$ and p-value between AUC drop and plaintext KL using k samples using SGD (left) and LBFGS (right).
Data Addition $\pi$ $r$ p-value
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$0.182 3.65e-02
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$0.162 6.27e-02
$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$0.184 3.47e-02
$\mathrm{KL}_{\mathcal{X}}$0.162 7.13e-02
Gender 0.097 2.65e-01
Race 0.018 8.29e-01
Age 0.053 5.33e-01
Table 2. Using the eICU dataset, we measure the Pearson correlation $r$ between the strategy $\pi$ and data addition AUC drop. p-values below .05.05.05.05 are bolded.
Appendix B Hyperparamter Tuning
We obtain Score(X,Y) by training a Logistic Regression model using SGD. We find that SGD requires hyper-parameter tuning in order to perform well when evaluated on Brier Score Loss. We used Optuna to perform hyper-parameters search. The hyper-parameters we use for plaintext scores are:
  1. (1)
    learning rate: $0.0795$
  2. (2)
    patience: 2
  3. (3)
    tolerance: 0.000117
  4. (4)
    momentum: 0.886
  5. (5)
    weight decay: 1.81e-9
  6. (6)
    dampening: .0545
The hyper-parameters for the encrypted model:
  1. (1)
    learning rate: $0.0974$
  2. (2)
    patience: 5
  3. (3)
    tolerance: 0.000132
  4. (4)
    momentum: 0.907
  5. (5)
    weight decay: 8.14e-7
  6. (6)
    dampening: .0545
Appendix C Correlations between Encrypted Scores and Plaintext Scores
Hospital $\rho(\mathrm{KL}_{\mathcal{X}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}})$ p-value $\rho(\mathrm{KL}_{\mathcal{X}\mathcal{Y}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}})$ p-value
73 0.945 1.118e-05 1.000 0.0
264 0.973 5.142e-07 0.945 1.118e-05
420 0.982 8.403e-08 0.991 3.763e-09
243 0.973 5.142e-07 0.909 1.056e-04
338 0.973 5.142e-07 0.982 8.403e-08
443 0.964 1.852e-06 0.882 3.302e-04
199 0.991 3.763e-09 0.973 5.142e-07
458 0.873 4.546e-04 0.964 1.852e-06
300 0.455 1.601e-01 0.691 1.857e-02
188 0.718 1.280e-02 0.864 6.117e-04
252 0.873 4.546e-04 0.809 2.559e-03
167 0.764 6.233e-03 0.891 2.335e-04
Table 3. Spearman Correlations $\rho$ for encrypted (in CrypTen) and plaintext (in PyTorch) KL-based methods
On Table 3, we measure the Spearman correlations between $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$, and between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ for all hospitals. We find that all hospitals have statistically significant correlations with the exception of hospital 300’s $\rho$($\mathrm{KL}_{\mathcal{X}}$,$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$)

### v2

@@
 Privacy-Preserving Dataset Combination
-Abstract.
+Abstract
 Access to diverse, high-quality datasets is crucial for machine learning model performance, yet data sharing remains limited by privacy concerns and competitive interests, particularly in regulated domains like healthcare.
@@
 We present SecureKL, a privacy-preserving framework that enables organizations to identify beneficial data partnerships without exposing sensitive information.
-Building on recent advances in dataset combination methods, we develop a secure multiparty computation protocol that maintains strong privacy guarantees while achieving ¿90% correlation with plaintext evaluations.
+Building on recent advances in dataset combination methods, we develop a secure multiparty computation protocol that maintains strong privacy guarantees while achieving $>90\%$ correlation with plaintext evaluations.
 In experiments with real-world hospital data, SecureKL successfully identifies beneficial data partnerships that improve model performance for intensive care unit mortality prediction while preserving data privacy.
@@
 Our code is publicly available at https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
-isbn: 978-1-4503-XXXX-X/18/06
-1. Introduction
-Empirical scaling laws have established clear relationships between model performance and three key factors: compute, data, and model size (Hestness2017DeepLS, ; Kaplan2020ScalingLF, ).
-These relationships have driven remarkable improvements across computer vision, language processing, speech recognition, reinforcement learning, and healthcare applications (sun2017revisiting, ; mahajan2018exploring, ; brown2020language, ; hoffmann2022training, ; park2019specaugment, ; zhang2020pushing, ; ashvin2020accelerating, ; reed2022generalist, ; sheller2020federated, ; mckinney2020international, ).
-Beyond simple scaling, the diversity of training data has proven crucial for enhancing model robustness to distribution shifts and mitigating performance disparities across demographic groups (shen2024data, ; miller2021accuracy, ).
+1 Introduction
+Empirical scaling laws have established clear relationships between model performance and three key factors: compute, data, and model size [1, 2].
+These relationships have driven remarkable improvements across computer vision, language processing, speech recognition, reinforcement learning, and healthcare applications [3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
+Beyond simple scaling, the diversity of training data has proven crucial for enhancing model robustness to distribution shifts and mitigating performance disparities across demographic groups [13, 14].
 However, access to data varies significantly across entities and domains.
 While large tech companies have the data and compute resources to train the foundational models that now dominate general tasks, smaller players often lack such access.
-Domain-specific data is also becoming increasingly valuable for fine-tuning these general models (alsentzer2019publicly, ; gururangan2020don, ; lee2020biobert, ), creating a competitive advantage for the data owners.
-As a result, entities with domain-specific data are more reluctant to share it for free, opting instead to sell it in emerging data markets (acemoglu2022too, ; huang2021toward, ; liang2018survey, ).
+Domain-specific data is also becoming increasingly valuable for fine-tuning these general models [15, 16, 17], creating a competitive advantage for the data owners.
+As a result, entities with domain-specific data are more reluctant to share it for free, opting instead to sell it in emerging data markets [18, 19, 20].
 This dynamic particularly disadvantages smaller organizations, which often lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
@@
 In healthcare, for example, patient data is heavily regulated by laws such as the Health Insurance Portability and Accountability Act (HIPAA), which imposes strict data-sharing constraints to protect patient privacy.
-Moreover, accumulating data from multiple sources introduces the risk of domain shift, where data from different distributions may degrade model performance instead of improving it (koh2021wilds, ; wang2022generalizing, ; taori2020measuring, ).
-Contrary to the intuition that more data always leads to better performance, combining datasets does not guarantee improvements — in fact, performance can decrease when incorporating additional data sources (bradley2021unrepresentative, ; meng2018statistical, ; shen2024data, ).
+Moreover, accumulating data from multiple sources introduces the risk of domain shift, where data from different distributions may degrade model performance instead of improving it [21, 22, 23].
+Contrary to the intuition that more data always leads to better performance, combining datasets does not guarantee improvements — in fact, performance can decrease when incorporating additional data sources [24, 25, 13].
 This non-monotonic behavior means that carefully selecting which datasets to combine is crucial, as using all available data may actually perform worse than using an optimal subset.
-While existing methods aim to identify datasets combinations that can improve performance, they often assume access to all relevant datasets (shen2024data, ; ilyas2022datamodels, ).
+While existing methods aim to identify datasets combinations that can improve performance, they often assume access to all relevant datasets [13, 26].
 However, this is impractical because entities are reluctant to share their data due to privacy risks or the competitive value they associate with it.
 This reluctance to share creates a bottleneck in improving model performance.
-In high-stakes settings like healthcare, this bottleneck is particularly detrimental, as access to diverse data has the potential to drive significant advancements in patient care and outcomes (beam2018big, ; ghassemi2019practical, ; chen2020treating, ; beaulieu2019trends, ).
+In high-stakes settings like healthcare, this bottleneck is particularly detrimental, as access to diverse data has the potential to drive significant advancements in patient care and outcomes [27, 28, 29, 30].
 To address these challenges, we propose SecureKL, a privacy-preserving approach for guiding dataset combinations without requiring direct data or model sharing.
 Our work makes three primary technical contributions:
-  1. (1)
+  1. 1.
     We introduce a framework for practical secure data combination that enables organizations to evaluate potential partnerships while maintaining data privacy.
 Our framework categorizes existing approaches by their privacy leakage risks and provides a systematic way to assess the trade-offs between data utility and privacy preservation.
-  2. (2)
-    We extend KL-XY score (shen2024data, ) with a secure multiparty protocol that enables privacy-preserving evaluation of potential data partnerships.
-Our protocol maintains strong privacy guarantees while utilizing the complete underlying datasets, achieving ¿90% correlation with plaintext computations.
-  3. (3)
+  2. 2.
+    We extend KL-XY score [13] with a secure multiparty protocol that enables privacy-preserving evaluation of potential data partnerships.
+Our protocol maintains strong privacy guarantees while utilizing the complete underlying datasets, achieving $>90\%$ correlation with plaintext computations.
+  3. 3.
     Through extensive evaluation in low-data, high-stakes settings, we demonstrate that our method successfully identifies beneficial data partnerships for intensive care unit (ICU) mortality prediction, improving classifier performance for the source hospital.
-  4. (4)
-    In simulated scenarios requiring selection of three partner hospitals, Private-KL-XY outperforms alternative selection strategies including demographic-based selection (using gender, race, and age), and limited-sample plaintext selection.
+  4. 4.
+    In experimental scenarios requiring selection of three partner hospitals, Private-KL-XY outperforms alternative selection strategies including demographic-based selection (using gender, race, and age), and limited-sample plaintext selection.
 We argue that our method presents an appealing trade-off of privacy and utility by preserving privacy for both parties while using all the underlying data.
-2. Problem Setup
+2 Problem Setup
 Consider a binary prediction task for ICU patient mortality based on electronic medical records.
@@
 Using only their data, $H_{o}$ can train a model $\mathcal{M}$ with parameters $\theta$ to achieve:
-(Baseline Performance) $\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$
+$\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$ (Baseline Performance)
 where $f$ is their chosen algorithm with parameter $\theta$.
@@
 By combining datasets, i.e., $\mathcal{D}_{T}=\{\mathcal{D}_{i}\mid i\in T\}\cup\mathcal{D}_{o}$, $H_{o}$ can potentially achieve better results:
-(Combined Performance) $\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$
+$\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$ (Combined Performance)
 We define the potential improvement from data addition as $\delta_{T}=\delta_{(o,T)}=\mathrm{AUC}_{T}-\mathrm{AUC}_{o}$.
@@
 Formally, given $n\leq N$, we seek a strategy $\pi$ that selects $n$ target datasets $T=\pi(\mathcal{D}_{o},n)$ to maximize model utility:
-(Ideal Dataset Combination) $\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset{[1...N]\choose n}}\mathrm{AUC}_{T}$
+$\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset{[1...N]\choose n}}\mathrm{AUC}_{T}$ (Ideal Dataset Combination)
 Practical Considerations.
@@
 Kullback–Leibler Divergence.
-Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work (shen2024data, ).
-KL divergence (kullback1951information, ), also called information gain (quinlan1986induction, ), describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
-(Kullback–Leibler Divergence) $\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$
-Because computing KL-divergence on datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial,  (shen2024data, ) proposes two groups of scores to make this divergence approximation tractable from small samples.
+Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work [13].
+KL divergence [31], also called information gain [32], describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
+$\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$ (Kullback–Leibler Divergence)
+Because computing KL-divergence on datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial,  [13] proposes two groups of scores to make this divergence approximation tractable from small samples.
 Specifically, score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates — with the goal of inferring dataset membership.
 Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
-(KL-XY Score) $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$
+$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$ (KL-XY Score)
 Details are described in Section 3.
@@
 This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation.
-This privacy preservation model incentivizes collaboration, improving upon methods in  (shen2024data, ) .
+This privacy preservation model incentivizes collaboration, improving upon methods in  [13] .
 MPC Preliminary
-To secure this divergence computation cryptographically, Secure Multiparty Computation (MPC) (yao1982protocols, ; shamir1979share, ) protocols are leveraged.
+To secure this divergence computation cryptographically, Secure Multiparty Computation (MPC) [33, 34] protocols are leveraged.
 Specifically, in $\mathrm{SecureKL}$, each party encodes $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ to preserve privacy for both parties.
-This is implemented with the research framework CrypTen (knott2021crypten, ), specialized for MPC and machine learning.
+This is implemented with the research framework CrypTen [35], specialized for MPC and machine learning.
 Our algorithmic and engineering details are in Sections  3 and  4, respectively.
@@
 In order to make privacy boundaries tractable, we make the following additional assumptions:
-  1. (1)
+  1. 1.
     Existing knowledge is not private.
@@
 Hospitals frequently know of each other’s resources, and the available ICU units are contentious, not kept secret.
-  2. (2)
+  2. 2.
     Uniformity of $|\mathcal{D}_{i}|$.
 Though each hospital gets to price their data and set their own budget, for generality, the uniformity assumption allows us to use the number of additional data sources $n$ as the main ”budget proxy” across different strategies.
-  3. (3)
+  3. 3.
     Legal risks of sharing any data are omnipresent in high stakes domains.
@@
 This abstraction side-steps legal discussion, which would go beyond the scope of our paper.
-  4. (4)
+  4. 4.
     No malice is assumed on any of the parties involved, as each hospital wants to authentically sell their data and set up a potential collaboration.
@@
 We note our limitations in Section 4.5.
-3. Methods
-Figure 1.
-Our Method $\mathrm{SecureKL}(\mathcal{D}_{o},\mathcal{D}_{i})$.
+3 Methods
+Figure 1: Our Method $\mathrm{SecureKL}(\mathcal{D}_{o},\mathcal{D}_{i})$.
 Each side encrypts their data, and a model is trained on their joint data.
 Using private KL-based measures, their distance is computed, and the final result is revealed after decryption, which requires both parties.
-[…]
-Data is encrypted on each side.
 We summarize our data acquisition strategies differentiated by leakage risks, which correlate with potential costs:
 Category 1, medium-to-high leakage, sharing raw data.
-$\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered (lds, ).
+$\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered [36].
 Though leakage can be controlled through $k$, the data is inherently sensitive.
-The underlying distance uses  (shen2024data, ) ’s KL-XY Score.
+The underlying distance uses  [13] ’s KL-XY Score.
 Category 2, medium leakage, sharing summary statistics.
@@
 Our method $\pi_{p}(n)$ selects data sources based on privacy-preserving measure for data combination, specifically Private KL-XY.
-3.1. Trivial Baseline: Blind Selection
+3.1 Trivial Baseline: Blind Selection
 Blind selection refers to the process when no information is provided.
@@
 This random strategy may evade selection biases and help gather diverse data.
-Yet, prior work  (shen2024data, ) suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
-3.2. Sharing Summary Statistics
+Yet, prior work  [13] suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
+3.2 Sharing Summary Statistics
 A relaxation to sharing no sensitive data is to share metadata.
@@
 Then, we use the distributional distance between $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, characterized by their $L_{2}$-distance through $\phi$:
-(Demographic-based Strategy) $\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$
-3.3. KL-based Methods, in Plaintext
+$\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$ (Demographic-based Strategy)
+3.3 KL-based Methods, in Plaintext
 $\pi_{s}(n,k)$ assumes each of the candidate hospitals will share a set of raw data.
@@
 Though leakage can be controlled through $k$, the data is inherently sensitive.
-This is implemented with KL-based measures similar to (shen2024data, ) .
+This is implemented with KL-based measures similar to [13] .
 To recap, $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a ”metric” that satisfies triangle inequality.
 Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be ”close” to the target distribution $P_{i}$, but not the other way around:
-(Ideal Estimator) $\mathrm{KL}(P_{o}||P_{i})=\int_{x\in\mathcal{X}}\log\frac{P_{o}(\mathop{}\!\mathrm{d}x)}{P_{i}(\mathop{}\!\mathrm{d}x)}P_{i}(\mathop{}\!\mathrm{d}x)$
+$\mathrm{KL}(P_{o}||P_{i})=\int_{x\in\mathcal{X}}\log\frac{P_{o}(\mathop{}\!\mathrm{d}x)}{P_{i}(\mathop{}\!\mathrm{d}x)}P_{i}(\mathop{}\!\mathrm{d}x)$ (Ideal Estimator)
 Because we only have access to finite data $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, approximations are needed.
@@
 Thus the joint distribution of features and labels from both the source and target are included, with the goal of deriving an efficient estimator for $\mathrm{KL}(P_{o}||P_{i})$ that captures distributional shift from source to target.
-Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model (Cox1958TheRA, ) on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covaraiates – with the goal of inferring dataset membership.
+Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model [37] on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covaraiates – with the goal of inferring dataset membership.
 A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful.
-(shen2024data, ) established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task.
+[13] established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task.
 We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
 Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
-(KL-X) $\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$
-(KL-XY) $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$
-We focus on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, and reproduce (shen2024data, ) ’s results that it is predictive of downstream change in AUC.
+$\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$ (KL-X)
+$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$ (KL-XY)
+We focus on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, and reproduce [13] ’s results that it is predictive of downstream change in AUC.
 Let the score function $g_{\mathrm{KL}}$ be the approximate of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$.
 The strategy selects the most likely hospital with the closest distance under the measure:
-(KL-based Strategy, in plaintext) $\pi_{s}(n=1,k=K)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{KL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$
+$\pi_{s}(n=1,k=K)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{KL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$ (KL-based Strategy, in plaintext)
 When only a subset is available, this function is adjusted by swapping $\mathcal{D}_{i}$ for $\mathcal{D}_{i}^{\prime}\subseteq\mathcal{D}_{i}$ where $|\mathcal{D}_{i}^{\prime}|=k$.
 We denote the full dataset size as $K=|\mathcal{D}_{i}|$.
-3.4. SecureKL: Private KL-based Method
+3.4 SecureKL: Private KL-based Method
 Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public).
-Specifically we leverage the MPC based framework provided by CrypTen (knott2021crypten, ), a library designed for privacy-preserving machine learning, to implement private $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
+Specifically we leverage the MPC based framework provided by CrypTen [35], a library designed for privacy-preserving machine learning, to implement private $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
 As illustrated in Figure 1, the logistic regression as well as the scoring need to be implemented in private.
@@
 Denote the private encoding of $x$ as $[x]$.
-(Secure KL-X) $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$
-(Secure KL-XY) $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}([x,y])).$
+$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$ (Secure KL-X)
+$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}([x,y])).$ (Secure KL-XY)
 Let the score function $g_{\mathrm{SKL}}$ be the secure approximation of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$.
 The strategy selects the most likely hospital with the closest distance under the measure:
-(SecureKL Strategy, encrypted) $\pi_{p}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{SKL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$
+$\pi_{p}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{SKL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$ (SecureKL Strategy, encrypted)
 As shown in Figure 1, any KL-based measure $g_{\mathrm{SKL}}$ can be adapted to our setup.
@@
 We note the engineering limitations in Section 4.5.3.
-4. Experiments
-4.1. Experimental Setup
-We validate our method and demonstrate its applicability using the eICU Collaborative Research Dataset (Pollard2018TheEC, ), which contains over $200,000$ admissions from $208$ hospitals across the United States.
-Following the data cleaning and exclusion criteria outlined by (Water2023YetAI, ) and (shen2024data, ), we selected the $12$ hospitals with the highest number of patient visits (each with at least $2000$ patients) as our $\mathbf{H}$.
+4 Experiments
+4.1 Experimental Setup
+We validate our method and demonstrate its applicability using the eICU Collaborative Research Dataset [38], which contains over $200,000$ admissions from $208$ hospitals across the United States.
+Following the data cleaning and exclusion criteria outlined by [39] and [13], we selected the $12$ hospitals with the highest number of patient visits (each with at least $2000$ patients) as our $\mathbf{H}$.
 Each strategy would compute with the same $K=3000$ records, as the total available data per hospital.
@@
 The strategy comparisons described in Section 3 are implemented using $1500$ samples and the AUC is evaluated on $400$ samples for all of our experiments unless otherwise noted.
-This follows training and evaluation protocols in Yet Another ICU Benchmark (Water2023YetAI, ).
-For the data combination experiments that compute AUC change $\delta_{i}$ or $\delta_{T}$, to match  (shen2024data, ) , we take $1500$ random samples from each selected dataset and combine it with $1500$ samples from $\mathcal{D}_{o}$ (fixed across all experiments).
+This follows training and evaluation protocols in Yet Another ICU Benchmark [39].
+For the data combination experiments that compute AUC change $\delta_{i}$ or $\delta_{T}$, to match  [13] , we take $1500$ random samples from each selected dataset and combine it with $1500$ samples from $\mathcal{D}_{o}$ (fixed across all experiments).
 Implementing $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be privacy-preserving requires training logistic regression model in private.
 This is used in the private setting to estimate $g_{\mathrm{SKL}}$$\mathrm{Score}([X])$ or $\mathrm{Score}([X,Y])$ – for each pair of hospitals.
-Our experiments train encrypted logistic regression in CrypTen (knott2021crypten, ) using the library’s SGD optimizer.
-To ensure a fair comparison between the scores obtained through plaintext and encrypted settings, we re-implement plaintext Score($X$) and Score($X,Y$) using logistic regression with SGD in PyTorch (paszke2019pytorch, ).
-This is because encrypted version of L-BFGS – the optimizer prior work (shen2024data, ) uses in plaintext-only with Scikit-Learn (scikit-learn, )– is not available in CrypTen, though it leads to better downstream performance.
-Hyper-parameter tuning for SGD in private and plain text are performed independently, with the details in Appendix B.
-4.2. Experimental Questions
+Our experiments train encrypted logistic regression in CrypTen [35] using the library’s SGD optimizer.
+To ensure a fair comparison between the scores obtained through plaintext and encrypted settings, we re-implement plaintext Score($X$) and Score($X,Y$) using logistic regression with SGD in PyTorch [40].
+This is because encrypted version of L-BFGS – the optimizer prior work [13] uses in plaintext-only with Scikit-Learn [41]– is not available in CrypTen, though it leads to better downstream performance.
+Hyper-parameter tuning for SGD in private and plain text are performed independently, with the details in Appendix 8.
+4.2 Experimental Questions
 We ask three sets of questions:
-  1. (1)
+  1. 1.
     Consistency: Does using multiparty implementation sacrifice original measure’s effectiveness?
@@
 In addition to computing Spearman’s rank correlation coefficient of the KL-scores, we probe the discrepancy of the downstream effect between $\{\delta_{i}|i=\pi_{s}(n=1,k=K)\}$ using plaintext $\mathrm{KL}_{\mathcal{X}},\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ with all the underlying data and $\{\delta_{i}|\pi=\pi_{p}\}$ using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ for each source hospital $H_{o}\in\mathbf{H}$.
-  2. (2)
+  2. 2.
     Positivity: Does our method pick hospitals that reliably improve performance?
@@
     Lastly, Section 4.4.1 compare with different strategies proposed in Section 3, and Section 4.4.2 analyze the benefits of using private dataset combination $\mathrm{SKL}$.
-  3. (3)
+  3. 3.
     Error analysis: If our privacy-preserving method is not the dominant strategy against alternatives including limited data accessibility, why is that?
@@
 Lastly, we note engineering hurdles to scale to the real-world in Section 4.5.3. We hereby detail our results.
-4.3. Consistency Between Plaintext and Encrypted Computations
+4.3 Consistency Between Plaintext and Encrypted Computations
 Our encrypted computations are programmed with CrypTen, when the plaintext counterpart is PyTorch-only.
@@
 Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$.
-After applying Hochberg false discovery rate correction (Benjamini1995ControllingTF, ), our p-values remain significant.
+After applying Hochberg false discovery rate correction [42], our p-values remain significant.
 This range is an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations.
 When we unify the SGD hyperparameters, we indeed get a tighter range.
-For all 12 hospitals, see appended Appendix C for details.
+For all 12 hospitals, see appended Appendix 9 for details.
 Downstream Model Improvements
@@
 Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of $-0.164$ with $\delta_{i}$.
-For all strategies’ correlations with ground truth at $n=1$, see Appendix A.
-Figure 2.
-AUC change $\delta_{T}$, after including top-$3$ hospital per strategy, over all source hospitals.
+For all strategies’ correlations with ground truth at $n=1$, see Appendix 7.
+Figure 2: AUC change $\delta_{T}$, after including top-$3$ hospital per strategy, over all source hospitals.
 Our private strategy $\pi_{p}$ is compared with (a) demographic-based $\pi_{d}$ for gender, age, and race.
 (b) plain text limited-sample $\pi_{s}(k=K)$ for $k=300$ (10%) and $k=30$ (1%).
-4.4. Positivity of SecureKL and Practical Implications
+4.4 Positivity of SecureKL and Practical Implications
 Overall Positivity
@@
 Overall, 34 out of the 36 dataset combinations we evaluate on have an AUC improvement $\delta_{T}>0$, suggesting that $\pi_{p}$ is a reasonable strategy for selecting hospital dataset combinations with a high expected return $\mathbb{E}[P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)]$ for the source hospital from using our strategy.
-4.4.1. Comparing With Alternative Strategies
+4.4.1 Comparing With Alternative Strategies
 Other strategies – $\pi_{0}$, $\pi_{d}$, and $\pi_{s}$ – can also arrive at positive datasets.
 Comparing private method $\pi_{p}(n=3)$ to other strategies at $n=3$, we find the following results, illustrated in Figure 2:
-  1. (1)
+  1. 1.
 […]
-  2. (2)
+  2. 2.
     Demographic-based strategies underperform compared to $\pi_{p}$ on average.
@@
 Similarly, $\pi_{d}$-age has a median of $0.014$, and $\pi_{d}$-race has a median of $0.008$, both lower than $\pi_{p}$’s median.
-  3. (3)
+  3. 3.
     Plaintext small-sample strategies, $\pi_{s}$, outperform all demographic-based methods but slightly underperform relative to $\pi_{p}$.
@@
 In contrast, demographic-based and plaintext small-sample strategies exhibit greater variability, with some strategies improving performance for specific subsets of hospitals but underperforming or degrading results in others.
-4.4.2. SecureKL Analysis
+4.4.2 SecureKL Analysis
 After establishing that $\pi_{p}$ with $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$is a robust strategy in practical downstream performance, we hereby synopsize the benefits of $\mathrm{SecureKL}$ and elaborate on their practical implications.
@@
 This strong guarantee can significantly ease the tension related to privacy and compliance in setting up a collaboration, leading to a practical ”data appraisal stage” in data-limited high stakes domains.
-In the case where that output can be sensitive, i.e., when hospitals query each other multiple times and accrue information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response (dwork2014algorithmic, ).
+In the case where that output can be sensitive, i.e., when hospitals query each other multiple times and accrue information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response [43].
 In theory, any data combination method (if Turing-complete) can be made private; yet, in practice, balancing the right trade-off of utility and privacy is non-trivial.
 Barring engineering difficulties, not all algorithms readily adapt efficiently in private.
-Prior work  (xu2022data, ) included the trained model and test data in private; while relatively exact, complex methods would exacerbate the same operational limitations discussed in Section 4.5.3.
+Prior work  [44] included the trained model and test data in private; while relatively exact, complex methods would exacerbate the same operational limitations discussed in Section 4.5.3.
 Gain from Data Availability.
@@
 This suggests a domain-specific alternative to sharing a large amount of data for some source hospital, and points to future directions to using secure computation on a minimal-sized sample dataset for minimal performance overhead while remaining private.
-4.5. Error Analysis
-4.5.1. Underlying Score Limitations
+4.5 Error Analysis
+4.5.1 Underlying Score Limitations
 Data addition algorithms underpin the effectiveness of our method.
@@
 This behavior stems from the underlying measure, not from adding secure computation: in Figure 4 and Figure 5, the encrypted performance closely follows that of plaintext performance, for both good and bad downstream correlations.
-4.5.2. Sometimes, Not All Underlying Data Is Needed
+4.5.2 Sometimes, Not All Underlying Data Is Needed
 Relatedly, when seeing a few samples can successfully identify useful candidate hospitals, $\pi_{p}$ does not always outperform $\pi_{s}$ on small samples.
@@
 We further note that our method still clearly applies to encrypted computation on a smaller data set under data minimization.
-Figure 3. Hospital 243: Underlying KL-score performs poorly
-[…]
-NEED UPDATING some conclusion and description
-Figure 4. Hospital 420: Underlying KL-score performs well
-[…]
-NEED UPDATING
-Figure 5.
-Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{p}(k=30)$ and $\pi_{0}$.
+Figure 3: Hospital 243: Underlying KL-score performs poorly
+Figure 4: Hospital 420: Underlying KL-score performs well
+Figure 5: Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{p}(k=30)$ and $\pi_{0}$.
 Middle: All strategies perform similarly.
@@
 In all panels, the bars represent the standard deviation.
-[…]
-NEED UPDATING
-4.5.3. Encrypted Computation Limitations
+4.5.3 Encrypted Computation Limitations
 Engineering a secure system for machine learning requires both machine learning and software engineering knowledge.
 We share our our code and method, but also note potential challenges with deploying our secure computation:
-  1. (1)
+  1. 1.
     Operational: engineering personnel limitations.
 While our implementation requires little cryptographic knowledge to deploy, it still needs technically-trained staff at each participating hospital to collaborate and maintain.
-  2. (2)
+  2. 2.
     Engineering: Extending any MPC protocol is non-trivial, as security engineering is a specialized skill.
 While $\mathrm{SecureKL}$ applies broadly to other underlying scores in multi-party setups, every new algorithm requires software engineering - prototyping, tuning, debugging —which can be especially costly for hospitals.
-  3. (3)
+  3. 3.
     Framework Limitation: While CrypTen is designed to accommodate PyTorch, it is a research tool where not all plain text functionalities are implemented.
 Writing optimizers – such as L-BFGS – and custom operators that are not readily available requires both machine learning and cryptography knowledge.
-  4. (4)
+  4. 4.
     Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly.
-As detailed in Appendix  B, our hyperparameters for SGD are indeed different in encrypted and plaintext settings.
+As detailed in Appendix  8, our hyperparameters for SGD are indeed different in encrypted and plaintext settings.
 However, encrypted computation hides loss curves and training details by default, complicating development.
-5. Related Work
-5.1. Data Valuation and Pricing
+5 Related Work
+5.1 Data Valuation and Pricing
 The question of how to assess the impact and therefore worth of data has been well studied.
 Data valuation as a field seeks to quantify the contribution of individual data points or datasets to model performance.
-Shapley value-based approaches provide theoretically grounded valuations but scale poorly to large datasets (ghorbani2019data, ).
-More efficient methods include influence functions (koh2017understanding, ) and leave-one-out testing (cook1979influential, ).
-Recent work has extended these concepts to dataset-level valuation (jia2019towards, ).
-As practitioners create larger datasets, the emergence of data marketplaces has sparked interest in data pricing mechanisms (kumar2020marketplace, ).
-Query-based pricing (koutris2015query, ) and outcome-driven valuations (radic2024pricing, ) aim to balance seller compensation with buyer utility.
+Shapley value-based approaches provide theoretically grounded valuations but scale poorly to large datasets [45].
+More efficient methods include influence functions [46] and leave-one-out testing [47].
+Recent work has extended these concepts to dataset-level valuation [48].
+As practitioners create larger datasets, the emergence of data marketplaces has sparked interest in data pricing mechanisms [49].
+Query-based pricing [50] and outcome-driven valuations [51] aim to balance seller compensation with buyer utility.
 While these approaches inform fair data exchange, they typically assume direct access to data, unlike our privacy-preserving method.
-5.2. Alternative Approaches to Data Sharing
+5.2 Alternative Approaches to Data Sharing
 Recent work has explored several approaches to mitigate data sharing constraints while maintaining model performance.
@@
 Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy.
-Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data (goncalves2020generation, ), medical imaging data (thambawita2022singan, ), and electronic health records (baowaly2019synthesizing, ).
+Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data [52], medical imaging data [53], and electronic health records [54].
 These methods can preserve statistical properties of the original data while providing differential privacy guarantees.
-Transforming data into a similar form that de-sensitizes certain attributes can be desirable (drechsler2011synthetic, ; howe2017synthetic, ; nikolenko2021synthetic, ; gonzales2023synthetic, ; sweeney2002k, ).
-Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself (jordon2021hide, ).
-Additionally, outside of the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks (narayanan2006break, ; jordon2021hide, ).
-However, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables (goncalves2020generation, ).
-For tabular data, methods like CTGAN (xu2019modeling, ) and TVAE (xu2019modeling, ) have demonstrated ability to learn complex distributions while preserving correlations between features.
-However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance (assefa2020generating, ).
-Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees (jordon2018pate, ).
-While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset (yang2024tabular, ).
+Transforming data into a similar form that de-sensitizes certain attributes can be desirable [55, 56, 57, 58, 59].
+Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself [60].
+Additionally, outside of the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks [61, 60].
+However, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables [52].
+For tabular data, methods like CTGAN [62] and TVAE [62] have demonstrated ability to learn complex distributions while preserving correlations between features.
+However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance [63].
+Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees [64].
+While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset [65].
 Public pre-training and private fine-tuning
 Transfer learning via public pretraining has become increasingly popular for domains with limited private data access.
-BioBERT (lee2020biobert, ) and ClinicalBERT (alsentzer2019publicly, ) demonstrated that pretraining on PubMed abstracts and clinical notes can improve performance on downstream medical tasks.
-Similar approaches have emerged in other regulated domains, including FinBERT (liu2021finbert, ) for financial applications.
+BioBERT [17] and ClinicalBERT [15] demonstrated that pretraining on PubMed abstracts and clinical notes can improve performance on downstream medical tasks.
+Similar approaches have emerged in other regulated domains, including FinBERT [66] for financial applications.
 However, the effectiveness of transfer learning depends heavily on domain alignment.
-One study showed that continued pretraining on domain-specific data significantly outperforms generic pretraining when domains differ substantially (gururangan2020don, ).
-This presents challenges for highly specialized fields where public data may not capture domain-specific patterns (le2022language, ).
+One study showed that continued pretraining on domain-specific data significantly outperforms generic pretraining when domains differ substantially [16].
+This presents challenges for highly specialized fields where public data may not capture domain-specific patterns [67].
 Recent work has explored methods to quantify and optimize domain adaptation.
-Adaptive pretraining strategies (howard2018universal, ) and domain-specific vocabulary augmentation (gururangan2020don, ) have shown promise in bridging domain gaps.
+Adaptive pretraining strategies [68] and domain-specific vocabulary augmentation [16] have shown promise in bridging domain gaps.
 However, these approaches still require substantial compute and may not fully capture specialized domain knowledge present in private datasets.
-5.3. Benefits of Data Scaling Beyond Performance
+5.3 Benefits of Data Scaling Beyond Performance
 Recent work has demonstrated that increasing dataset size and diversity yields benefits beyond raw performance metrics.
-Large-scale training data has been shown to improve model robustness to distribution shifts (taori2020measuring, ) and reduce demographic performance disparities (chen2018my, ).
-Studies of vision models trained on increasingly large datasets show improved out-of-distribution generalization (miller2021accuracy, ).
-Similarly, language models trained on diverse data demonstrate better performance across different domains and demographic groups (dodge2021documenting, ).
-5.4. Secure and Confidential Computation
+Large-scale training data has been shown to improve model robustness to distribution shifts [23] and reduce demographic performance disparities [69].
+Studies of vision models trained on increasingly large datasets show improved out-of-distribution generalization [14].
+Similarly, language models trained on diverse data demonstrate better performance across different domains and demographic groups [70].
+5.4 Secure and Confidential Computation
 Secure and confidential computation encompasses cryptographic techniques that protect information privacy during computation.
@@
 While it represents the gold standard for encrypted computation, adapting it to modern machine learning is challenging due to computational constraints from growing ciphertext size.
-FHE implementations typically use lattice-based schemes requiring periodic “bootstrapping” (key refreshing and noise reduction through re-encryption) via methods like the CKKS scheme (cheon2017homomorphic, ).
+FHE implementations typically use lattice-based schemes requiring periodic “bootstrapping” (key refreshing and noise reduction through re-encryption) via methods like the CKKS scheme [71].
 This introduces cryptographic parameters that non-experts struggle to configure effectively.
 Secure Multi-party Computation (S-MPC)
-SMPC enables multiple parties to compute functions over private inputs while revealing only the final output (yao1982protocols, ; shamir1979share, ).
-An MPC system uses key exchanges, encryption schemes, and secure communication to ensure only encrypted data leaves owner control (micali1987play, ).
+SMPC enables multiple parties to compute functions over private inputs while revealing only the final output [33, 34].
+An MPC system uses key exchanges, encryption schemes, and secure communication to ensure only encrypted data leaves owner control [72].
 Though generally faster than FHE, SMPC’s engineering complexity and communication overhead can limit adoption.
 Traditional private training approaches that completely hide data can also impede essential model development tasks like inspection, monitoring, and debugging.
-5.5. Secure Data Combination
+5.5 Secure Data Combination
 Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance.
-Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data (aono2017privacy, ).
-However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets (mcmahan2017communication, ).
+Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data [73].
+However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets [74].
 More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment.
-These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets (leung2019towards, ; chakraborty2024privacy, ).
-Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information (duan2021flexible, ).
+These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets [75, 76].
+Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information [77].
 Several systems have been developed to facilitate secure data combination in specific domains.
-In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance (raisaro2018m, ).
-Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality (liu2021finbert, ).
-5.6. Other Privacy-Preserving Methods
+In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance [78].
+Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality [66].
+5.6 Other Privacy-Preserving Methods
 Federated Learning.
-Cross-silo federated, decentralized, and collaborative machine learning  (mcmahan2017communication, ; li2020federated, ; bonawitztowards, ; kairouz2021advances, ) focus on acquiring more data through improved data governance and efficient system design.
-Healthcare machine learning is considered especially suitable, as health records are often isolated (rieke2020future, ; xu2021federated, ; nguyen2022federated, ).
+Cross-silo federated, decentralized, and collaborative machine learning  [74, 79, 80, 81] focus on acquiring more data through improved data governance and efficient system design.
+Healthcare machine learning is considered especially suitable, as health records are often isolated [82, 83, 84].
 Yet, even though no raw data is shared, model parameters or gradients flow through the system.
-As the federated computing paradigm offer no privacy guarantee, the system is vulnerable to model inversion (geiping2020inverting, ) and gradients leakage attacks (boenisch2023curious, ; zhu2019deep, ).
-A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimisation is traded off with privacy (lyu2022privacy, ; raynal2024conflict, ).
+As the federated computing paradigm offer no privacy guarantee, the system is vulnerable to model inversion [85] and gradients leakage attacks [86, 87].
+A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimisation is traded off with privacy [88, 89].
 Building on the insight that useful data is often disparately owned, we tackle the specific incentive problem between pairs of data players where one side trains the model, instead of scaling up a federation (number of parties) to address data access issues.
 We thus focusing on making this exchange efficient, accurate, and private.
-Compared to vanilla Federated Learning, an MPC system (shamir1979share, ; yao1982protocols, ; bonawitz2017practical, ; knott2021crypten, ) provides stronger guarantee in terms of input security.
+Compared to vanilla Federated Learning, an MPC system [34, 33, 90, 35] provides stronger guarantee in terms of input security.
 Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements.
-Our work extends the line of works  (xu2022data, ; yang2024fedfed, ; bonawitz2017practical, ) that demonstrates the potential of incorporating MPC in various federated scenarios.
-On the practical side, unlike mobile-based networks for secure federated learning protocols (bonawitztowards, ), our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
+Our work extends the line of works  [44, 91, 90] that demonstrates the potential of incorporating MPC in various federated scenarios.
+On the practical side, unlike mobile-based networks for secure federated learning protocols [80], our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
 Differential Privacy
-Differential privacy (DP) (dwork2006differential, ) offers formal privacy guarantees for sharing data and training machine learning models.
+Differential privacy (DP) [92] offers formal privacy guarantees for sharing data and training machine learning models.
 While DP mechanisms can protect individual privacy when releasing model outputs or aggregated statistics, they face significant limitations for inter-organizational data sharing.
-The primary challenge is that DP operates on already-pooled data, but organizations are often unwilling to share their raw data in the first place (dwork2014algorithmic, ).
+The primary challenge is that DP operates on already-pooled data, but organizations are often unwilling to share their raw data in the first place [43].
 Even when organizations are willing to share data, the privacy guarantees of DP come at a substantial cost to utility, particularly in machine learning applications.
-DP-SGD, the standard approach for training deep neural networks with differential privacy, significantly degrades model performance compared to non-private training (abadi2016deep, ).
-This performance impact is especially pronounced in data-constrained settings, where recent work has shown that large models rely heavily on memorization of rare examples that DP mechanisms tend to obscure (feldman2020neural, ).
+DP-SGD, the standard approach for training deep neural networks with differential privacy, significantly degrades model performance compared to non-private training [93].
+This performance impact is especially pronounced in data-constrained settings, where recent work has shown that large models rely heavily on memorization of rare examples that DP mechanisms tend to obscure [94].
 The privacy-utility trade-off becomes even more challenging when dealing with high-dimensional data or complex learning tasks.
-Studies have demonstrated that achieving meaningful privacy guarantees while maintaining acceptable model performance requires prohibitively large datasets (bagdasaryan2020backdoor, ).
-This limitation is particularly problematic in specialized domains like healthcare, where data is inherently limited and performance requirements are stringent (geyer2017differentially, ).
-Recent work has attempted to improve the privacy-utility trade-off through advanced composition theorems and adaptive privacy budget allocation (papernot2021tempered, ).
-However, these approaches still struggle to match the performance of non-private training, especially when working with modern deep learning architectures (tramer2020differentially, ).
+Studies have demonstrated that achieving meaningful privacy guarantees while maintaining acceptable model performance requires prohibitively large datasets [95].
+This limitation is particularly problematic in specialized domains like healthcare, where data is inherently limited and performance requirements are stringent [96].
+Recent work has attempted to improve the privacy-utility trade-off through advanced composition theorems and adaptive privacy budget allocation [97].
+However, these approaches still struggle to match the performance of non-private training, especially when working with modern deep learning architectures [98].
 While differential privacy offers important theoretical guarantees, our work focuses on the practical challenge of enabling data owners to evaluate potential partnerships before sharing any data, addressing a key barrier to collaboration that DP alone cannot solve.
-6. Conclusion
+6 Conclusion
 Our work demonstrates that privacy-preserving data valuation can help organizations identify beneficial data partnerships while maintaining data sovereignty.
@@
 Future work should explore extending these techniques to handle more complex data types and dynamic distribution shifts while maintaining strong privacy guarantees.
-Appendix A Correlation with downstream performance
-On Table 1, we report the Pearson correlations between $\pi_{s}(k=K)$ for $k\in\{3,30,300,3000\}$ and $\delta_{i}$.
-On Table 2, we report the Pearson correlations between different strategies and $\delta_{i}$.
-k $\rho$ p-value
-3 -0.063 4.70e-01
-30 -0.082 3.47e-01
-300 -0.059 5.00e-01
-3000 -0.184 3.47e-02
-k $\rho$ p-value
-3 -0.158 7.02e-02
-30 0.167 5.60e-02
-300 -0.097 2.70e-01
-3000 -0.284 9.47e-04
-Table 1. $\rho$ and p-value between AUC drop and plaintext KL using k samples using SGD (left) and LBFGS (right).
-Data Addition $\pi$ $r$ p-value
-$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$0.182 3.65e-02
-$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$0.162 6.27e-02
-$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$0.184 3.47e-02
-$\mathrm{KL}_{\mathcal{X}}$0.162 7.13e-02
-Gender 0.097 2.65e-01
-Race 0.018 8.29e-01
-Age 0.053 5.33e-01
-Table 2. Using the eICU dataset, we measure the Pearson correlation $r$ between the strategy $\pi$ and data addition AUC drop. p-values below .05.05.05.05 are bolded.
-Appendix B Hyperparamter Tuning
-We obtain Score(X,Y) by training a Logistic Regression model using SGD.
-We find that SGD requires hyper-parameter tuning in order to perform well when evaluated on Brier Score Loss.
-We used Optuna to perform hyper-parameters search.
-The hyper-parameters we use for plaintext scores are:
-  1. (1)
-    learning rate: $0.0795$
-  2. (2)
-    patience: 2
-  3. (3)
-    tolerance: 0.000117
-  4. (4)
-    momentum: 0.886
-  5. (5)
-    weight decay: 1.81e-9
-  6. (6)
-    dampening: .0545
-The hyper-parameters for the encrypted model:
-  1. (1)
-    learning rate: $0.0974$
-  2. (2)
-    patience: 5
-  3. (3)
-    tolerance: 0.000132
-  4. (4)
-    momentum: 0.907
-  5. (5)
-    weight decay: 8.14e-7
-  6. (6)
-    dampening: .0545
-Appendix C Correlations between Encrypted Scores and Plaintext Scores
-Hospital $\rho(\mathrm{KL}_{\mathcal{X}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}})$ p-value $\rho(\mathrm{KL}_{\mathcal{X}\mathcal{Y}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}})$ p-value
-73 0.945 1.118e-05 1.000 0.0
-264 0.973 5.142e-07 0.945 1.118e-05
-420 0.982 8.403e-08 0.991 3.763e-09
-243 0.973 5.142e-07 0.909 1.056e-04
-338 0.973 5.142e-07 0.982 8.403e-08
-443 0.964 1.852e-06 0.882 3.302e-04
-199 0.991 3.763e-09 0.973 5.142e-07
-458 0.873 4.546e-04 0.964 1.852e-06
-300 0.455 1.601e-01 0.691 1.857e-02
-188 0.718 1.280e-02 0.864 6.117e-04
-252 0.873 4.546e-04 0.809 2.559e-03
-167 0.764 6.233e-03 0.891 2.335e-04
-Table 3. Spearman Correlations $\rho$ for encrypted (in CrypTen) and plaintext (in PyTorch) KL-based methods
-On Table 3, we measure the Spearman correlations between $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$, and between $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ for all hospitals.
-We find that all hospitals have statistically significant correlations with the exception of hospital 300’s $\rho$($\mathrm{KL}_{\mathcal{X}}$,$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$)

### v3

@@
 Abstract
-Access to diverse, high-quality datasets is crucial for machine learning model performance, yet data sharing remains limited by privacy concerns and competitive interests, particularly in regulated domains like healthcare.
-This dynamic especially disadvantages smaller organizations that lack resources to purchase data or negotiate favorable sharing agreements.
-We present SecureKL, a privacy-preserving framework that enables organizations to identify beneficial data partnerships without exposing sensitive information.
-Building on recent advances in dataset combination methods, we develop a secure multiparty computation protocol that maintains strong privacy guarantees while achieving $>90\%$ correlation with plaintext evaluations.
-In experiments with real-world hospital data, SecureKL successfully identifies beneficial data partnerships that improve model performance for intensive care unit mortality prediction while preserving data privacy.
-Our framework provides a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages.
-These results demonstrate the potential for privacy-preserving data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data resources.
-Our code is publicly available at https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
+Privacy concerns and competitive interests impede data access for machine learning, due to the inability to privately assess external data’s utility.
+This dynamic disadvantages smaller organizations that lack resources to aggressively pursue data-sharing agreements.
+In data-limited scenarios, not all external data is beneficial, and collaborations suffer especially in heavily regulated domains: metrics that aim to assess external data given a source e.g., approximating their KL-divergence, require accessing samples from both entities pre-collaboration, hence violating privacy (Compton et al., 2023; Shen et al., 2024).
+This conundrum disempowers legitimate data-sharing, leading to a false “privacy-utility trade-off”.
+To resolve privacy and uncertainty tensions simultaneously, we introduce SecureKL, the first secure protocol for dataset-to-dataset evaluations with zero privacy leakage, designed to be applied preceding data sharing.
+SecureKL evaluates a source dataset against candidates, performing dataset divergence metrics internally with private computations, all without assuming downstream models.
+On real-world data, SecureKL achieves high consistency ($90\%$ correlation with non-private counterparts) and successfully identifies beneficial data collaborations in highly-heterogeneous domains (ICU mortality prediction across hospitals and income prediction across states).
+Our results highlight that secure computation maximizes data utilization, outperforming privacy-agnostic utility assessments that leak information.
 1 Introduction
-Empirical scaling laws have established clear relationships between model performance and three key factors: compute, data, and model size [1, 2].
-These relationships have driven remarkable improvements across computer vision, language processing, speech recognition, reinforcement learning, and healthcare applications [3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
-Beyond simple scaling, the diversity of training data has proven crucial for enhancing model robustness to distribution shifts and mitigating performance disparities across demographic groups [13, 14].
-However, access to data varies significantly across entities and domains.
-While large tech companies have the data and compute resources to train the foundational models that now dominate general tasks, smaller players often lack such access.
-Domain-specific data is also becoming increasingly valuable for fine-tuning these general models [15, 16, 17], creating a competitive advantage for the data owners.
-As a result, entities with domain-specific data are more reluctant to share it for free, opting instead to sell it in emerging data markets [18, 19, 20].
-This dynamic particularly disadvantages smaller organizations, which often lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
-Scaling datasets in regulated domains is particularly challenging due to legal constraints and unpredictable outcomes from altering the training data composition.
-In healthcare, for example, patient data is heavily regulated by laws such as the Health Insurance Portability and Accountability Act (HIPAA), which imposes strict data-sharing constraints to protect patient privacy.
-Moreover, accumulating data from multiple sources introduces the risk of domain shift, where data from different distributions may degrade model performance instead of improving it [21, 22, 23].
-Contrary to the intuition that more data always leads to better performance, combining datasets does not guarantee improvements — in fact, performance can decrease when incorporating additional data sources [24, 25, 13].
-This non-monotonic behavior means that carefully selecting which datasets to combine is crucial, as using all available data may actually perform worse than using an optimal subset.
-While existing methods aim to identify datasets combinations that can improve performance, they often assume access to all relevant datasets [13, 26].
-However, this is impractical because entities are reluctant to share their data due to privacy risks or the competitive value they associate with it.
-This reluctance to share creates a bottleneck in improving model performance.
-In high-stakes settings like healthcare, this bottleneck is particularly detrimental, as access to diverse data has the potential to drive significant advancements in patient care and outcomes [27, 28, 29, 30].
-To address these challenges, we propose SecureKL, a privacy-preserving approach for guiding dataset combinations without requiring direct data or model sharing.
-Our work makes three primary technical contributions:
-  1. 1.
-    We introduce a framework for practical secure data combination that enables organizations to evaluate potential partnerships while maintaining data privacy.
-Our framework categorizes existing approaches by their privacy leakage risks and provides a systematic way to assess the trade-offs between data utility and privacy preservation.
-  2. 2.
-    We extend KL-XY score [13] with a secure multiparty protocol that enables privacy-preserving evaluation of potential data partnerships.
-Our protocol maintains strong privacy guarantees while utilizing the complete underlying datasets, achieving $>90\%$ correlation with plaintext computations.
-  3. 3.
-    Through extensive evaluation in low-data, high-stakes settings, we demonstrate that our method successfully identifies beneficial data partnerships for intensive care unit (ICU) mortality prediction, improving classifier performance for the source hospital.
-  4. 4.
-    In experimental scenarios requiring selection of three partner hospitals, Private-KL-XY outperforms alternative selection strategies including demographic-based selection (using gender, race, and age), and limited-sample plaintext selection.
-We argue that our method presents an appealing trade-off of privacy and utility by preserving privacy for both parties while using all the underlying data.
+Figure 1: Privacy can dis-incentivize data collaborations.
+Without seeing external data, an organization has two strategies: i.
+blind default $\pi_{0}$: randomly selecting partnerships causes hesitation and hinders partnerships.
+ii.
+private evaluation $\pi_{p}$: securely assessing datasets before commitment.
+Since Hestness et al.
+(2017), empirical works have predetermined data as a key driver to performance gains, via the so-called “scaling laws” (Kaplan et al., 2020; Brown et al., 2020).
+Yet, accessing and combining datasets is persistently challenging.
+As datasets have evolved from small to larger (Sun et al., 2017), more diverse (Zhang et al., 2020), and more compute-optimal (Hoffmann et al., 2022), the field of machine learning continues to seek more data  (Mahajan et al., 2018; Brown et al., 2020; Park et al., 2019; Reed et al., 2022; Sheller et al., 2020) and better ways to combine it  (Shen et al., 2023; Nguyen et al., 2023; Petty et al., 2024; Li et al., 2024; McKinney et al., 2020).
+Strategically combining data from different sources promises enhanced models, but disempowers smaller organizations.
+While diverse, high quality data often improves performance, robustness, and fairness (Miller et al., 2021), access to such data significantly varies across entities and domains (Tenopir et al., 2011; L’Heureux et al., 2017).
+As domain-specific data becomes increasingly valuable (Alsentzer et al., 2019; Gururangan et al., 2020; Lee et al., 2020), data-owning entities are more reluctant to share it for free, opting instead to sell it in emerging data markets (Acemoglu et al., 2022; Huang et al., 2021; Liang et al., 2018).
+This dynamic disproportionately handicaps smaller organizations who lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
+Organizations may hesitate to commit to a potential partnership when unsure about the benefits.
+As Figure 1 illustrates, this “commitment issue” is not solely a privacy issue; it’s the inability to privately assess an external dataset’s utility before partnerships.
+This evaluation is, however, nontrivial.
+For example, adding in-domain data does not necessarily result in more performant models.
+Due to the precarious nature of domain shifts, machine learning models are sensitive to additional training data, often unpredictably when the source dataset is small (e.g., a single hospital’s data) – a phenomenon known as “the dataset combination problem” (Koh et al., 2021; Wang et al., 2022; Taori et al., 2020; Bradley et al., 2021; Meng, 2018; Shen et al., 2024).
+Figure 2: The Dataset Combination Problem.
+Real-world data collaborations are inherently uncertain.
+AUC change for a source entity, after incorporating external data across hospitals and states.
+Each entry representas training on a pairwise data combination for a test entity (x-axis) Each column represents training on a data combination for a fixed test dataset, formed by pairing the fixed entity with an external dataset in the y-axis.
+Left: In eICU (Pollard et al., 2018), 10 out of 12 hospitals may see their mortality prediction model degrade for some potential hospital partners.
+Right: In Folktables (Ding et al., 2021), combining with random state leads to worse income prediction in 10 out of 15 states.
+(red is bad; exact values are reported in Appendix 15)
+Ideally, all target data should be considered to reduce uncertainty in costly data collaborations.
+Yet, datasets owned by separate entities cannot be directly and fully accessed, significantly limiting the practicality of non-private dataset measures (Shen et al., 2024; Ilyas et al., 2022).
+Our work directly addresses this crux by recognizing both privacy and competitive incentives.
+First, before committing to acquiring unseen data, we enable organizations to privately gauge the relative utility of candidate datasets.
+Second, we provide strong privacy guarantees required of entities operating under stringent regulations e.g.
+healthcare providers to navigate data acquisition.
+Developers are often uncertain about the most effective model before more data becomes available.
+This renders a secure data appraisal stage by Xu et al.
+(2022), which requires model parameters, not applicable.
+In this opportunistic setting, we ask:
+Can we ascertain the differential utility of prospective datasets, without knowing the final model?
+We introduce SecureKL, the first divergence-based dataset measure with zero privacy leakage.
+Our key insight is that private divergence computations (via secure multiparty computation (MPC)) are more data-efficient than sharing samples, thus achieving high performance.
+By preserving performance, SecureKL presents a compelling guarantee: for both parties, privacy is fully protected while data utilization is maximized.
+Contributions
+A novel secure dataset-to-dataset evaluation protocol SecureKL (SKL) that reduces uncertainty in data utility under limited data and budget, producing privacy-preserving measures while using the maximum available samples.
+SKL achieves a $90\%$ correlation with privacy-violating counterparts across two real-world heterogeneous domains.
+Empirically, on ICU mortality prediction, SKL reliably chooses beneficial hospital(s) to partner with, outperforming data-leaking alternatives, including using demographic summaries or sharing data subsets.
+Impact
+We provide a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages.
+Our major advantage lies in reliability, especially when small organizations cannot afford to invest in detrimental partnerships.
+These results demonstrate the potential for wider data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data.
+Our code is available, and can be readily deployed to demonstrate potential data value preceding collaborations [^1] https://github.com/kere-nel/secure-data-eval.
+2 Background
+2.1 Dataset Combination Problem
+In high-stakes domains, incorporating additional datasets may degrade the model.
+In healthcare scenarios, both Compton et al.
+(2023) and Shen et al.
+(2024) showed that blindly acquiring new datasets can degrade model performance.
+We extend these findings by replicating the effect under our own experimental setup.
+As shown in Figure 2, combining data from one hospital or state with an external source often leads to inconsistent, and sometimes negative, changes in model performance.
+This non-monotonic behavior highlights the need to evaluate data partners before embarking on a full-fledged collaboration.
+2.2 Trading Off Sample Utility For “Privacy”
+Figure 3: Non-private evaluation strategies.
+iii.
+sub sampling $\pi_{s}$: a subset of the target’s data is shared.
+iv.
+demographic summaries $\pi_{d}$: the target entity discloses distributions by protected attributes, i.e.
+age, gender, or race.
+Before an agreement, informed negotiations become impossible when entities do not expose data.
+In practice, data owners frequently resort to tiny samples (Centers for Medicare & Medicaid Services, 2025) or summary statistics (e.g., race, gender, age) for making decisions (Figure 3).
+Yet, in data-limited settings, model performance is highly sensitive to new input.
+These heuristics are fickle, as sparse traits or limited samples often fail to capture the entire dataset’s nuance and complexity, especially in heterogeneous domains, creating a perceived privacy-utility trade-off.
+We will show that secure computation, on the other hand, can avoid this tradeoff by measuring over entire datasets while maintaining privacy.
 3 Problem Setup
@@
 Using only their data, $H_{o}$ can train a model $\mathcal{M}$ with parameters $\theta$ to achieve:
-$\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$ (Baseline Performance)
+$\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$ (Baseline AUC)
 where $f$ is their chosen algorithm with parameter $\theta$.
@@
 By combining datasets, i.e., $\mathcal{D}_{T}=\{\mathcal{D}_{i}\mid i\in T\}\cup\mathcal{D}_{o}$, $H_{o}$ can potentially achieve better results:
-$\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$ (Combined Performance)
+$\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$ (Combined AUC)
 We define the potential improvement from data addition as $\delta_{T}=\delta_{(o,T)}=\mathrm{AUC}_{T}-\mathrm{AUC}_{o}$.
 To add a single additional data source by setting $T=\{i\}$, the improvement is $\delta_{i}=\delta_{(o,i)}=\mathrm{AUC}_{i}-\mathrm{AUC}_{o}$.
-This leads to our central question: Without seeing target data, how does a hospital ascertain potential data sources to combine with?
+This leads to our central question:
+Without seeing target data, how does a hospital ascertain potential data sources to combine with?
 Formally, given $n\leq N$, we seek a strategy $\pi$ that selects $n$ target datasets $T=\pi(\mathcal{D}_{o},n)$ to maximize model utility:
-$\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset{[1...N]\choose n}}\mathrm{AUC}_{T}$ (Ideal Dataset Combination)
+$\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset\binom{[1...N]}{n}}\mathrm{AUC}_{T}$ (Ideal Combination)
 Practical Considerations.
-Computing every subset $T\subset{[1...N]\choose n}$’s associated $\delta_{T}$ is exponential in $n$.
+Computing every subset $T\subset\binom{[1...N]}{n}$’s associated $\delta_{T}$ is exponential in $n$.
 To make this problem tractable, we make two key assumptions.
@@
 With the ultimate objective of improving the source hospital’s prediction task, we fix $H_{o}$; to compare the trade-offs between strategies in Section 4, we apply each $\pi$ greedily to select top-$n$ institution(s) for $H_{o}$ without replacement.
-Second, in in data constrained settings, we aim to maximize the probability of positive improvement: $P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)$.
+Second, in data constrained settings, we aim to maximize the probability of positive improvement: $P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)$.
 Kullback–Leibler Divergence.
-Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work [13].
-KL divergence [31], also called information gain [32], describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
-$\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$ (Kullback–Leibler Divergence)
-Because computing KL-divergence on datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial,  [13] proposes two groups of scores to make this divergence approximation tractable from small samples.
-Specifically, score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates — with the goal of inferring dataset membership.
-Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
-$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$ (KL-XY Score)
-Details are described in Section 3.
-Privacy Model for $\pi_{p}$.
+Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work (Shen et al., 2024).
+KL divergence (Kullback and Leibler, 1951), also called information gain (Quinlan, 1986), describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
+$\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$ (KL-Divergence)
+Because computing KL-divergence on hospital datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial due to the high dimensionality of the data, Shen et al.
+(2024) proposes two groups of scores to make this divergence approximation tractable from small samples.
+We adopt one such score, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, which trains a classifier to distinguish between source and target samples and uses the classifier’s output probabilities to estimate dataset divergence.
+We describe the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score in detail in Section 4.2.
+Privacy Model
 We operate under a semi-honest privacy model—also known as honest-but-curious or passive security—where parties follow protocols but may probe intermediate values.
@@
 This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation.
-This privacy preservation model incentivizes collaboration, improving upon methods in  [13] .
-MPC Preliminary
-To secure this divergence computation cryptographically, Secure Multiparty Computation (MPC) [33, 34] protocols are leveraged.
-Specifically, in $\mathrm{SecureKL}$, each party encodes $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ to preserve privacy for both parties.
-This is implemented with the research framework CrypTen [35], specialized for MPC and machine learning.
-Our algorithmic and engineering details are in Sections  3 and  4, respectively.
-For related secure techniques, see Section 5.4.
-Additional Assumptions
-Generally, we consider high stakes domains where disparate data may have additive benefits to the existing data.
-In order to make privacy boundaries tractable, we make the following additional assumptions:
+This privacy preservation model incentivizes collaboration, improving upon Shen et al.
+(2024) and  Ilyas et al.
+(2022).
+Dataset Divergence and Downstream Utility
+Building on the finding of Miller et al.
+(2021) where a model’s in-distribution performance is related to its out-of-distribution performance (across all model choices), we confer that dataset divergence does predict downstream model’s performance after combining the data (Table 3), using metrics introduced in Shen et al.
+(2024) (Section 4.2).
+Intuitively, in-distribution quality is paramount in low-data settings, where dataset divergence can capture greater complexity and nuance than accessing a few traits.Yet, privacy is unresolved: divergence measures entail accessing both entities’ data (Shen et al., 2024; Ilyas et al., 2022), posing significant risks for heavily-regulated entities who are liable for any data exposure (Centers for Medicare & Medicaid Services, 1996; Gervasi et al., 2022).
+Secure Multiparty Computation (MPC)
+To cryptographically secure our divergence computation, we use Secure Multiparty Computation (MPC)  (Yao, 1982; Shamir, 1979).
+MPC lets two or more parties to compute a function over private inputs, revealing only the final output (Goldwasser and Micali, 2019).
+Engineering machine learning workflows in private faces differs drastically from non-private machine learning engineering, due to 1.
+precision configurations and normalizations (to avoid blow-ups and underflows), 2.
+error control and performance engineering for machine learning in bit-limited spaces, and 3.
+debugging, hyperparameter tuning.
+Nevertheless, they are provably secure.
+Despite their non-trivial engineering, MPC programs enjoy strong security guarantees and relative ease of deployment.
+Even small organizations can deploy MPC without any specialized hardware.
+Thus, the algorithms developed and shared in SecureKL readily enable zero-leakage dataset-to-dataset evaluations before sharing data.
+Assumptions
+Consider high stakes domains where disparate data may have additive benefits to the existing data. To make privacy boundaries tractable, we make the following assumptions:
   1. 1.
@@
     Legal risks of sharing any data are omnipresent in high stakes domains.
-The risks with sharing sensitive data in $\pi_{d}$ and $\pi_{s}$ are not made explicit, but assumed to be ”medium” and ”medium-to-high” respectively.
+The risks with sharing sensitive data in data-leaking strategies, which we coin as $\pi_{d}$ (demographic distance) and $\pi_{s}$ (small sample), are not made explicit, but assumed to be ”moderate” and ”high” respectively.
 This abstraction side-steps legal discussion, which would go beyond the scope of our paper.
@@
 4 Methods
-Figure 1: Our Method $\mathrm{SecureKL}(\mathcal{D}_{o},\mathcal{D}_{i})$.
-Each side encrypts their data, and a model is trained on their joint data.
-Using private KL-based measures, their distance is computed, and the final result is revealed after decryption, which requires both parties.
-We summarize our data acquisition strategies differentiated by leakage risks, which correlate with potential costs:
-Category 1, medium-to-high leakage, sharing raw data.
-$\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered [36].
+Strategy Sub-strategies Leakage
+$\pi_{0}$ Blind (baseline) n/a zero
+$\pi_{p}$ Private (SecureKL) n/a minimal
+$\pi_{d}$ Demographic sex, gender, race moderate
+$\pi_{s}$ Subset Sampling 1%, 10%, 100% high
+Table 1: Concrete selection strategies $\pi$, differentiated by leakage risks. For a strategy $p$, a set of targets is chosen, i.e., $T\leftarrow\pi(\mathcal{D}_{o},n)$.
+4.1 Defining Dataset Acquisition Strategies
+Given the private and non-private data acquisition strategies illustrated in Figures 1 and 3, we associate privacy leakage as a primary cost of data partnerships.
+This section formalizes them by their leakage risks, summarized in Table 1.
+A, high leakage, sharing raw data.
+$\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered (Centers for Medicare & Medicaid Services, 2025).
 Though leakage can be controlled through $k$, the data is inherently sensitive.
-The underlying distance uses  [13] ’s KL-XY Score.
-Category 2, medium leakage, sharing summary statistics.
+The underlying distance metric follows the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score introduced by Shen et al.
+(2024) , which we define in the following subsection.
+B, moderate leakage, sharing summary statistics.
 $\pi_{d}(n)$ uses demographic metadata to guide data selection.
 This is implemented through ratio distance between source and target distributions, which may be considered aggregates therefore potentially not sensitive, such as when the underlying aggregation function $\phi$ is differentially private.
-Category 3, zero leakage, sharing no additional information besides what is assumed public.
+C, zero/minimal leakage, sharing no additional information besides what is assumed public, and what our method outputs, such as a score or a ranking.
 There are two methods: a.
@@
 b.
-Our method $\pi_{p}(n)$ selects data sources based on privacy-preserving measure for data combination, specifically Private KL-XY.
-3.1 Trivial Baseline: Blind Selection
-Blind selection refers to the process when no information is provided.
-$\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out.
-This random strategy may evade selection biases and help gather diverse data.
-Yet, prior work  [13] suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
-3.2 Sharing Summary Statistics
-A relaxation to sharing no sensitive data is to share metadata.
-While demographic traits are often causal and available, their exact cause in relation to the task is not a priori established (without a highly effective model), therefore their success in distributional-matching is not guaranteed to be strong.
-Additionally, in practice, the most effective model that results from data combination may or may not be causally-sound.
-Nevertheless, we posit alternative strategy $\pi_{d}(n)$ to find the demographically close candidates to guide data selection: Let $\phi:\mathcal{D}\to\mathbb{R}^{m}$ be an $m$-dimensional summary statistic of a demographic feature i.e.
-the racial distribution of patients.
-Then, we use the distributional distance between $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, characterized by their $L_{2}$-distance through $\phi$:
-$\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$ (Demographic-based Strategy)
-3.3 KL-based Methods, in Plaintext
-$\pi_{s}(n,k)$ assumes each of the candidate hospitals will share a set of raw data.
-In ICU data, simulate that a default of $1\%$ is shared, so $k=3000\times 1\%=30$, though we run experiments with $k\in\{3,30.300,3000\}$ (Section 4.3).
-Though leakage can be controlled through $k$, the data is inherently sensitive.
-This is implemented with KL-based measures similar to [13] .
-To recap, $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a ”metric” that satisfies triangle inequality.
-Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be ”close” to the target distribution $P_{i}$, but not the other way around:
+Our method $\pi_{p}(n)$ selects data sources based on privacy-preserving measure for data combination, specifically Private $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
+4.2 KL-based Methods, Without Privacy
+In subset-sampling strategy $\pi_{s}(n,k)$, each of the candidate entities will leak a set of raw data.
+$\pi_{s}$ is implemented with KL-based measures similar to Shen et al.
+(2024).
+Recall $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a “metric” that satisfies triangle inequality.
+Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be “close” to the target distribution $P_{i}$, but not the other way around:
 $\mathrm{KL}(P_{o}||P_{i})=\int_{x\in\mathcal{X}}\log\frac{P_{o}(\mathop{}\!\mathrm{d}x)}{P_{i}(\mathop{}\!\mathrm{d}x)}P_{i}(\mathop{}\!\mathrm{d}x)$ (Ideal Estimator)
@@
 Thus the joint distribution of features and labels from both the source and target are included, with the goal of deriving an efficient estimator for $\mathrm{KL}(P_{o}||P_{i})$ that captures distributional shift from source to target.
-Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model [37] on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covaraiates – with the goal of inferring dataset membership.
+Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model (Cox, 1958) on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates.
+The goal is to infer dataset membership using “proxy labels”, defined as follows:
+$\mathcal{I}(x,y)=\begin{cases}1&\text{if }(x,y)\in\mathcal{D}_{o}\\ 0&\text{if }(x,y)\in\mathcal{D}_{i}\end{cases}$
+A binary predictor is fit on this combined dataset, predicting $\mathcal{I}$ from $(\mathcal{X},\mathcal{Y})$ using logistic regression.
+The model’s output $p(x,y)$, also called the probability score, is Score(x,y)(x,y).
 A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful.
-[13] established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task.
+Shen et al.
+(2024) established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task, similar to Miller et al.
+(2021).
 We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
 Then, the resulting model’s probability score function $\text{Score}(\cdot)$: $\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
-$\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$ (KL-X)
 $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$ (KL-XY)
-We focus on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, and reproduce [13] ’s results that it is predictive of downstream change in AUC.
+We reproduce Shen et al.
+(2024) ’s results that $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ is predictive of downstream change in AUC in hospital setting (Table 4).
 Let the score function $g_{\mathrm{KL}}$ be the approximate of $\mathrm{KL}(\mathcal{D}_{o}||\mathcal{D}_{i})$.
@@
 We denote the full dataset size as $K=|\mathcal{D}_{i}|$.
+Though leakage can be controlled through $k$, yet the data is inherently sensitive.
+In ICU data, simulate that a default of $1\%$ is shared, so $k=3000\times 1\%=30$, though we run experiments with $k\in\{3,30,300,3000\}$.
 4.3 SecureKL: Private KL-based Method
+Figure 4: Our method SecureKL.
+Each side encrypts their data.
+Then, a model is privately trained on their joint data.
+Afterwards, their divergence is computed.
+Finally, only the final result of this dataset-to-dataset evaluation is revealed.
 Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public).
-Specifically we leverage the MPC based framework provided by CrypTen [35], a library designed for privacy-preserving machine learning, to implement private $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
+Specifically we leverage private tensors and secure gradient descent in CrypTen (Knott et al., 2021) to implement private $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
 As illustrated in Figure 4, the logistic regression as well as the scoring need to be implemented in private.
-Our code is publicly available  [^1] https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
 Denote the private encoding of $x$ as $[x]$.
-$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$ (Secure KL-X)
 $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}([x,y])).$ (Secure KL-XY)
@@
 Additionally, even though our implementation measures distance of data between one source and one target party, the setup readily extends to accommodating multiple parties.
-We note the engineering limitations in Section 4.5.3.
+Section 7.2 discusses potential deployment challenges.
+Figure 5: SecureKL: Overall Correctness.
+Rank correlation between SecureKL output and ground truth AUC change, $\delta_{i}$, from acquiring 11 additional dataset for a given source hospital $H_{o}$.
+We propose selecting data partner ranked by our secure system under $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score to reliably increase AUC gains.
+($|\mathbf{H}|=12$ hospitals; colored by source.)
+4.4 Trivially “Private” Baseline: Blind Selection
+$\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out.
+This random strategy may evade selection biases and help gather diverse data.
+Yet, prior work  (Shen et al., 2024) suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
+4.5 Alternative Data-leaking Baseline: Sharing Summary Statistics
+A relaxation to sharing no sensitive data is to share metadata.
+While demographic traits are often causal and available, their exact cause in relation to the task is not a priori established (without a highly effective model), therefore their success in distributional-matching is not guaranteed to be strong.
+Additionally, in practice, the most effective model that results from data combination may or may not be causally-sound.
+Nevertheless, we posit alternative strategy $\pi_{d}(n)$ to find the demographically close candidates to guide data selection: Let $\phi:\mathcal{D}\to\mathbb{R}^{m}$ be an $m$-dimensional summary statistic of a demographic feature i.e.
+the racial distribution of patients.
+Then, we use the distributional distance between $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$, characterized by their $L_{2}$-distance through $\phi$:
+$\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$ (Demographic-based Strategy)
 5 Experiments
-4.1 Experimental Setup
-We validate our method and demonstrate its applicability using the eICU Collaborative Research Dataset [38], which contains over $200,000$ admissions from $208$ hospitals across the United States.
-Following the data cleaning and exclusion criteria outlined by [39] and [13], we selected the $12$ hospitals with the highest number of patient visits (each with at least $2000$ patients) as our $\mathbf{H}$.
-Each strategy would compute with the same $K=3000$ records, as the total available data per hospital.
-We simulate the problem setup for each hospital with the $24$-hour mortality prediction task.
-The strategy comparisons described in Section 3 are implemented using $1500$ samples and the AUC is evaluated on $400$ samples for all of our experiments unless otherwise noted.
-This follows training and evaluation protocols in Yet Another ICU Benchmark [39].
-For the data combination experiments that compute AUC change $\delta_{i}$ or $\delta_{T}$, to match  [13] , we take $1500$ random samples from each selected dataset and combine it with $1500$ samples from $\mathcal{D}_{o}$ (fixed across all experiments).
-Implementing $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be privacy-preserving requires training logistic regression model in private.
-This is used in the private setting to estimate $g_{\mathrm{SKL}}$$\mathrm{Score}([X])$ or $\mathrm{Score}([X,Y])$ – for each pair of hospitals.
-Our experiments train encrypted logistic regression in CrypTen [35] using the library’s SGD optimizer.
-To ensure a fair comparison between the scores obtained through plaintext and encrypted settings, we re-implement plaintext Score($X$) and Score($X,Y$) using logistic regression with SGD in PyTorch [40].
-This is because encrypted version of L-BFGS – the optimizer prior work [13] uses in plaintext-only with Scikit-Learn [41]– is not available in CrypTen, though it leads to better downstream performance.
-Hyper-parameter tuning for SGD in private and plain text are performed independently, with the details in Appendix 8.
 5.1 Experimental Questions
-We ask three sets of questions:
   1. 1.
-    Consistency: Does using multiparty implementation sacrifice original measure’s effectiveness?
-Practically, we evaluate this through the analysis of private and plaintext scores.
-For our selected metrics, we expect AUC change to be negatively with KL-based measures—meaning the closer the additional target dataset is to the source hospital, the more the AUC will improve compared to other potential target datasets.
-Section 4.3 tests the correlation of our private scores and plaintext scores with full access (setting $k=K$).
-In addition to computing Spearman’s rank correlation coefficient of the KL-scores, we probe the discrepancy of the downstream effect between $\{\delta_{i}|i=\pi_{s}(n=1,k=K)\}$ using plaintext $\mathrm{KL}_{\mathcal{X}},\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ with all the underlying data and $\{\delta_{i}|\pi=\pi_{p}\}$ using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}},\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ for each source hospital $H_{o}\in\mathbf{H}$.
+    Consistency: Does MPC degrade the original measure’s effectiveness?
+Since MPC implementations introduce approximations, we first validate $\mathrm{SKL}$’s correctness by examining the correlation of private scores (SecureKL) vs.
+plaintext scores ($\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, with full data access).
+Then, for the consistency of choosing encryption in the hospital domains, we examine plaintext and encrypted versions’ correlation with downstream ground truth rankings across hospitals.
   2. 2.
-    Positivity: Does our method pick hospitals that reliably improve performance?
-If source dataset $D_{o}$ can only add data from $n$ more hospitals, does our measure lead to eventual AUC improvements?
-In Section 4.4, we test our framework on a multi-dataset combination experiment and find that it successful improves the source hospital’s downstream outcome.
-Specifically, when selecting a single additional data source ($n=1$), all but 2 hospitals improves, and when selecting top 2 or 3, all hospitals see a positive AUC change.
-This shows a consistent added benefit.
-    Lastly, Section 4.4.1 compare with different strategies proposed in Section 3, and Section 4.4.2 analyze the benefits of using private dataset combination $\mathrm{SKL}$.
+    Positivity: Does our method pick entities that reliably improve performance?
+If source dataset $D_{o}$ can only add data from $n$ more sources, does our measure lead to eventual AUC improvements?
+Specifically, in multi-dataset combination, we examine whether using $\mathrm{SKL}$ can improve the source hospital’s downstream outcome.
+When selecting a single (or a few) additional data source, how many hospitals improve with our method?
+Additionally, we compare our method with alternative, privacy-leaking strategies.
   3. 3.
-    Error analysis: If our privacy-preserving method is not the dominant strategy against alternatives including limited data accessibility, why is that?
-Section 4.5 performs additional analyses on (a) hospitals with low $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ correlations, and (b) hospitals lagging AUC improvements using the random strategy $\pi_{0}$ or limited-sample strategy $\pi_{s}$ for selecting $n=3$ candidate hospitals.
-Lastly, we note engineering hurdles to scale to the real-world in Section 4.5.3. We hereby detail our results.
+    Error analysis: If our privacy-preserving method is not the dominant strategy against alternatives including limited data accessibility, why?
+As we know, small and uncertain improvements for downstream tasks underscore the inherent difficulty of evaluating data utility without seeing the full data.
+We are especially interested in analyzing (a) hospitals with low $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ correlations, and (b) hospitals lagging AUC improvements using the random strategy $\pi_{0}$ or limited-sample strategy $\pi_{s}$.
+5.2 MPC implementation
+$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ includes training logistic regression model in private.
+We implement measures based on dataset divergence by building a custom logistic regression model over encrypted data leveraging CrypTen.
+The model parameters and input are encoded as 16-bit MPCTensors, ensuring that all computations, including forward passes, sigmoid activations, and gradient descent updates, are performed in private.
+Additional baseline details
+We additionally run our experiments on plaintext methods used in Shen et al.
+(2024), including the $\mathrm{KL}_{\mathcal{X}}$ measure, which is similar to $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ without using each data source’s labels:
+$\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$ (KL-X)
+To compare against $\mathrm{KL}_{\mathcal{X}}$ for baseline, we additionally implemented its encrypted version, $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$, as a $g_{\mathrm{SKL}}$ candidate.
+$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$ (Secure KL-X)
+Optimizers
+Because L-BFGS – the optimizer prior work (Shen et al., 2024) used in plaintext-only with Scikit-Learn (Pedregosa et al., 2011) – is not available as an encrypted version, our MPC experiments are facilitated with SGD optimizer.
+A fair comparison between the scores obtained through plaintext and encrypted settings necessitates re-implementing plaintext scores, Score($X$) and Score($X,Y$), using logistic regression with SGD in PyTorch (Paszke et al., 2019).
+The hyper-parameter tuning for SGD in private and plain text are performed independently, as they do not transfer.
+Appendix 12 will discuss hyperparameter-tuning in detail.
+5.3 Data and Model Setup
+eICU dataset.
+The downstream task is $24$-hour mortality prediction from ICU data using the eICU Collaborative Research Dataset (Pollard et al., 2018).
+This dataset contains >200,000>200,000 real-world admission records from $208$ hospitals across the United States.
+A note on data filtering.
+Because medical research is inherently complex, ensuring reproducibility on statistical methods can be significantly challenging.
+Water et al.
+(2023) proposes that the research community follow a shared set of tasks with fixed preprocessing pipelines that are clinically informed – essentially a machine learning training and evaluation protocol on eICU – in order to facilitate method verification on the same benchmark.
+Therefore, our work follows their data cleaning criteria and the evaluation protocol.
+Additionally, we use the hospital exclusion criteria in (Shen et al., 2024) to obtain top $12$ hospitals, where the most patient visits are collected (each with >2000>2000 patients).
+Downstream model and baselines for eICU.
+Recall that each strategy uses the same $K$ number of records per hospital – in our experiment, $K=3000$.
+For $\pi_{s}$ which leaks a subset $k$ of all samples, a default $k=1\%|\mathcal{D}_{i}|$ randomly drawn samples are shared.
+In ICU data, we run experiments on $\{0.1\%,1\%,10\%,100\%\}$.
+Following holistic benchmarking tools in  (Water et al., 2023), our strategy comparisons take $1500$ samples and the downstream model performance – $\mathrm{AUC}_{o}$, $\mathrm{AUC}_{T}$ – uses $400$ samples (unless otherwise noted).
+Specifically, the AUC change, $\delta_{i}$ or $\delta_{T}$, comes from 1.
+combining $1500$ random samples from each selected dataset and 2.
+combining it with $1500$ samples from $\mathcal{D}_{o}$, and 3.
+subtracting the baseline model’s AUC [^2] The samples are fixed across all experiments, the sample numbers are chosen to match (Shen et al., 2024)’s setup..
+Folktables dataset Though we primarily focus on hospital domain, we additionally validate using Folktables (Ding et al., 2021), predicting across 15 states an individual’s annual income exceeds \$50,000333Because the Folktables dataset contains simpler features and is low-dimensional, KL divergence is directly computed rather than estimated.
+For implementation details, see Appendix 15..
+The details of our processing, which diverges from that of eICU, is included in Appendix 15.
+6 Results and Analysis
+We organize our results around three research questions: whether using multiparty implementation sacrifices $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$’s efficacy (consistency; Section 6.1), whether our method reliably picks hospitals that improve performance (positivity; Section 6.2), and where our method may fail (error analysis and limitations; Section 6.3).
 6.1 Consistency Between Plaintext and Encrypted Computations
-Our encrypted computations are programmed with CrypTen, when the plaintext counterpart is PyTorch-only.
-We show using $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ lead to highly comparable behavior as $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}}$, respectively.
+Because our encrypted computation is the first implementation of dataset divergence in MPC, we ought to show that $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ lead to highly comparable behaviors as $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ and $\mathrm{KL}_{\mathcal{X}}$.
 Spearman’s Rank Correlation Coefficient for Underlying Scores
@@
 Between $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$, $\mathbb{E}_{H_{o}\sim\mathbf{H}}[\rho]=0.9303$ with a range of $[0.455,0.991]$, with 11 of 12 hospitals achieving p-values below $0.05$.
-After applying Hochberg false discovery rate correction [42], our p-values remain significant.
-This range is an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations.
-When we unify the SGD hyperparameters, we indeed get a tighter range.
+After applying Hochberg false discovery rate correction (Benjamini and Hochberg, 1995), our p-values remain significant.
+This range may be an artifact of sweeping hyperparameters independently in plaintext and encrypted optimisations, because sharing the same SGD hyperparameters would result in a tighter range.
 For all 12 hospitals, see appended Appendix 13 for details.
-Downstream Model Improvements
-We further simulate the effect by adding encryption through its impact on the downstream AUC.
-This examines whether there will be a shift in the full hospital ranking, if we switch from a plaintext setup to encrypted.
+Impact of Adding Security on AUC Correlation
+We further examine the effect by adding encryption through its impact on the downstream AUC, using how AUC improvements are ranked.
+This rank is compared with how secure measures (i.e., $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) and plaintext measures (i.e.
+$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) rank hospitals.
+This comparison investigates the extent of the shift in the full hospital ranking, when we switch from a plaintext setup to encrypted.
 For $H_{o}\sim\mathbf{H}$, we measure $\delta_{i}$ that results from adding $\mathcal{D}_{i}$ to $\mathcal{D}_{o}$ for all $i$.
 This correlates all target hospitals $\{H_{i}\}$ with their ground truths $\{\delta_{i}\}$ in the case of picking a single target hospital.
-We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.182$ and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be $-0.184$ ($99\%$ matching).
-Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of $-0.164$ with $\delta_{i}$.
+We find the linear coefficient for encrypted $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be 0.182-0.182 and plaintext $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be 0.184-0.184 ($99\%$ matching).
+Both $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}}$ have a linear coefficient of 0.164-0.164 with $\delta_{i}$.
 For all strategies’ correlations with ground truth at $n=1$, see Appendix 11.
-Figure 2: AUC change $\delta_{T}$, after including top-$3$ hospital per strategy, over all source hospitals.
-Our private strategy $\pi_{p}$ is compared with (a) demographic-based $\pi_{d}$ for gender, age, and race.
-(b) plain text limited-sample $\pi_{s}(k=K)$ for $k=300$ (10%) and $k=30$ (1%).
-4.4 Positivity of SecureKL and Practical Implications
-Overall Positivity
-We evaluate the practical utility of SecureKL by applying it in a multi-source data combination experiment, where $n\in\{1,2,3\}$.
+6.2 Positivity in Realistic Setup
+We apply SecureKL in a pragmatic multi-source data combination problem, where each strategy acquires $n$ datasets for $n\in\{1,2,3\}$.
+Overall Positivity.
 For $n=1$, we find that $\pi_{p}$ improves AUC in 10 out of the 12 hospitals.
@@
 Overall, 34 out of the 36 dataset combinations we evaluate on have an AUC improvement $\delta_{T}>0$, suggesting that $\pi_{p}$ is a reasonable strategy for selecting hospital dataset combinations with a high expected return $\mathbb{E}[P_{H_{o}\sim\mathbf{H}}(\delta_{T}>0)]$ for the source hospital from using our strategy.
-4.4.1 Comparing With Alternative Strategies
-Other strategies – $\pi_{0}$, $\pi_{d}$, and $\pi_{s}$ – can also arrive at positive datasets.
-Comparing private method $\pi_{p}(n=3)$ to other strategies at $n=3$, we find the following results, illustrated in Figure 2:
+Demographic Summary Strategies Subset-sampling Strategies Secure Strategy (Ours)
+Dataset $n$ $\pi_{d}$-gender $\pi_{d}$-age $\pi_{d}$-race $\pi_{s}(k=300)$ $\pi_{s}(k=30)$ $\pi_{p}\ \mathrm{Secure}\mathrm{KL}_{\mathcal{XY}}$
+eICU 1 $\textbf{0.020}\pm 0.023$ $012\pm 0.016 0.0$$14\pm 0.015 0.0$$12\pm 0.014 0.0$$\pm 0.017 0.011±0.0$$1\pm 0.018$
+2 $016\pm 0.016 0.0$$17\pm 0.017 0.0$$15\pm 0.013 0.0$$24\pm 0.020 0.0$$17\pm 0.019 0.0$$2\textbf{0.027}\pm$ 0.022
+[…]
+[…]
+2 $\textbf{0.015}\pm 0.019$ $\textbf{0.015}\pm 0.014$ $0\pm 0.013 0.011±0.0$$1\pm 0.014$ $014\pm 0.018 0.0$$010\pm 0.017$
+[…]
+Table 2: AUC improvements in mean and standard deviation, across all source regions for each strategy $\pi$, for eICU and Folktables setups.
+$n$ denotes the number of candidate datasets added to the source dataset.
+The small gains and high variance from adding selected datasets highlight the precarious nature of assessing data value in the real world.
+Bold indicates highest AUC improvement per $n$.
+Note: Only $\pi_{p}\ \mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ is private.
+Figure 6: AUC change $\delta_{T}$ over all strategies in eICU prediction (higher is better).
+Our private dataset evaluation strategy $\pi_{p}$ outperforms demographic-based strategy $\pi_{d}$ (left), and subset-sampling strategy $\pi_{s}$ for $k=300$ (10%) and $k=30$ (1%) (right), after combining source data with the top 33 candidates.
+Figure 7: AUC change $\delta_{T}$ over all strategies in Folktables dataset prediction (higher is better).
+All strategies exhibit comparable distributions, after combining data from top 3 candidates.
+In a noisy domain, our method is stable: it neither excels nor penalizes against non-private strategies.
+Comparing With Alternative Strategies
+Other strategies – $\pi_{0}$, $\pi_{d}$, and $\pi_{s}$ – can also arrive at “positive” datasets.
+Comparing private method to other strategies at $n=3$, i.e, $\pi_{p}(n=3)$, we describe our results in Figure 7:
   1. 1.
-[…]
+    $\pi_{p}$ (our method based on $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$) has a median $\delta_{T}$ of $0.020$, and a standard deviation of $0.015$.
+Our results indicate that for 5050% of the hospitals, $\pi_{p}$ gives a $\delta_{T}>=.02$.
+Compared to other strategies, $\pi_{p}$ has the highest median, the lowest standard deviation, and it is one of two strategies that improves performance for all hospitals.
   2. 2.
@@
 Compared to other strategies, it has the largest standard deviation ($0.024$), and it degrades the performance for some hospitals.
-In summary, our method $\pi_{p}$ achieves the highest AUC improvement on average with the lowest standard deviation, demonstrating consistent improvement for all hospitals.
-In contrast, demographic-based and plaintext small-sample strategies exhibit greater variability, with some strategies improving performance for specific subsets of hospitals but underperforming or degrading results in others.
-4.4.2 SecureKL Analysis
-After establishing that $\pi_{p}$ with $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$is a robust strategy in practical downstream performance, we hereby synopsize the benefits of $\mathrm{SecureKL}$ and elaborate on their practical implications.
-A Principled Approach To Data Minimization.
-Our major contribution is to match plaintext performance with no data sharing.
-Using MPC provides input privacy, meaning that if both hospitals only want to know the resulting score, the computation can be done without leaking original data.
-This strong guarantee can significantly ease the tension related to privacy and compliance in setting up a collaboration, leading to a practical ”data appraisal stage” in data-limited high stakes domains.
-In the case where that output can be sensitive, i.e., when hospitals query each other multiple times and accrue information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response [43].
-In theory, any data combination method (if Turing-complete) can be made private; yet, in practice, balancing the right trade-off of utility and privacy is non-trivial.
-Barring engineering difficulties, not all algorithms readily adapt efficiently in private.
-Prior work  [44] included the trained model and test data in private; while relatively exact, complex methods would exacerbate the same operational limitations discussed in Section 4.5.3.
-Gain from Data Availability.
-In contrast to limited-sample approaches, a key advantage for our method $\pi_{p}$ is that it takes advantage of all of the underlying data – generally impossible with non-secure methods for private data in heavily regulated domains.
-The general intuition is that data is localized; therefore, once a good target hospital is identified, we should acquire all of the data.
-It may be tempting to assert that we prefer the highest $k$ for data addition algorithms as well.
-In our experiments, while this is generally true, the smaller $k$ sometimes outperform larger $k$ in plaintext strategy $\pi_{s}$, which we investigate in Section 4.5 and in Figure 5.
-This occasionally non-monotonic behavior mirrors the challenge of data combination itself: even within one source dataset for the same estimator, more data is not necessarily better.
-This suggests a domain-specific alternative to sharing a large amount of data for some source hospital, and points to future directions to using secure computation on a minimal-sized sample dataset for minimal performance overhead while remaining private.
-4.5 Error Analysis
-4.5.1 Underlying Score Limitations
+In summary, our method $\pi_{p}$ achieves the highest AUC improvement on average with the lowest standard deviation, demonstrating more consistent improvements for all hospitals.
+While the average improvement of $\pi_{p}$ is small, demographic-based and plaintext small-sample strategies exhibit greater variability, with some strategies improving performance for specific subsets of hospitals but underperforming or degrading results in others.
+6.3 SecureKL Error Analysis
+Underlying Data Limitations
+In high-stakes domains, data partnerships are expensive, but potentially detrimental – this forms a challenging landscape for evaluating methods on real-world data.
+Indeed, as shown in Table 2, the AUC gain is small across all strategies, and the variance is high.
+This suggests that 3 hospitals’ data is likely still too small for the general task to the robust explains limited AUC gains, highlighting the need to maximize samples for informative decisions.
+The key distinction, however, is that privacy-leaking methods (demographic, small sample) and blind baseline risk performance declines in many hospitals while SKL consistently improves downstream tasks more reliably than alternatives, across all hospitals, over multiple runs.
+Figure 8: The correlation between AUC change $\delta_{i}$ and underlying score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ affects $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$’s efficacy.
+In Hospital 420, the $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ scores identify beneficial data ($\rho<0$).
+For Hospital 243, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ anti-correlates ($\rho>0$).
+Underlying Score Limitations
 Data addition algorithms underpin the effectiveness of our method.
@@
 This behavior stems from the underlying measure, not from adding secure computation: in Figure 8 and Figure 9, the encrypted performance closely follows that of plaintext performance, for both good and bad downstream correlations.
-4.5.2 Sometimes, Not All Underlying Data Is Needed
-Relatedly, when seeing a few samples can successfully identify useful candidate hospitals, $\pi_{p}$ does not always outperform $\pi_{s}$ on small samples.
+Sometimes, Not All Underlying Data Is Needed
+Relatedly, when seeing a few samples can successfully identify useful candidate hospitals, $\pi_{s}$ (which is on small samples) outperforms $\pi_{p}$ (which is on full samples).
 In the right panel of Figure 9, hospital 199, the smaller sample sizes achieve a score that better reflects ground truth as a data addition strategy.
@@
 We further note that our method still clearly applies to encrypted computation on a smaller data set under data minimization.
-Figure 3: Hospital 243: Underlying KL-score performs poorly
-Figure 4: Hospital 420: Underlying KL-score performs well
-Figure 5: Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{p}(k=30)$ and $\pi_{0}$.
-Middle: All strategies perform similarly.
-Right: $\pi_{p}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
-In all panels, the bars represent the standard deviation.
-4.5.3 Encrypted Computation Limitations
-Engineering a secure system for machine learning requires both machine learning and software engineering knowledge.
-We share our our code and method, but also note potential challenges with deploying our secure computation:
+7 Discussion
+7.1 SecureKL Contribution
+After establishing that $\pi_{p}$ with $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$is a robust strategy in practical downstream performance, we hereby summarize the benefits of $\mathrm{SecureKL}$ and elaborate on their practical implications.
+Matching Plaintext Performance in Downstream Tasks.
+Our major contribution is to match plaintext performance with no data sharing.
+Using MPC provides input privacy, meaning that if both hospitals only want to know the resulting score, the computation can be done without leaking original data.
+This strong guarantee can significantly ease the tension related to privacy and compliance in setting up a collaboration, leading to a practical ”data appraisal stage” in data-limited high stakes domains.
+Gain from Data Availability.
+In contrast to limited-sample approaches, a key advantage for our method $\pi_{p}$ is that it takes advantage of all the underlying data – generally impossible with non-secure methods for private data in heavily regulated domains.
+The general intuition is that data is localized; therefore, once a good target hospital is identified, we would prefer to acquire all of the data.
+It may be tempting to assert that we prefer the highest $k$ for data addition algorithms as well.
+In our experiments, while this is generally true, the smaller $k$ sometimes outperform larger $k$ in plaintext strategy $\pi_{s}$, which we investigate in Section 6.3 and in Figure 9.
+This occasionally non-monotonic behavior mirrors the challenge of data combination itself: even within one source dataset for the same estimator, more data is not necessarily better.
+This suggests the potential for a hospital-specific alternative to sharing a large amount of data for some source hospital, and points to future directions to using secure computation on a minimal-sized sample dataset for minimal performance overhead while remaining private.
+Potential Improvements to SecureKL In the case where that output can be sensitive, i.e., when a source hospital queries a target hospital multiple times and accrues information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response (Dwork et al., 2014).
+7.2 Potential Challenges to Broader Adoption
+Our code is readily usable by small organizations.
+While our approach generalizes to model-based measures (by substituting $g_{\mathrm{SKL}}$) and scales to multiple parties, our work also uncovered deployment limitations.
   1. 1.
@@
 While our implementation requires little cryptographic knowledge to deploy, it still needs technically-trained staff at each participating hospital to collaborate and maintain.
+This skill is similar to using pre-packaged software, cleaning data, and setting up network calls.
   2. 2.
-    Engineering: Extending any MPC protocol is non-trivial, as security engineering is a specialized skill.
-While $\mathrm{SecureKL}$ applies broadly to other underlying scores in multi-party setups, every new algorithm requires software engineering - prototyping, tuning, debugging —which can be especially costly for hospitals.
+    Engineering Extensions: Extending any MPC protocol is non-trivial, as security engineering is a specialized skill.
+While $\mathrm{SecureKL}$ applies broadly to other underlying scores in multi-party setups, validating a new MPC algorithm requires software engineering – prototyping, tuning, debugging – and numerical verification – akin to data analytics and research - likely requiring technical talents who can be especially costly for hospitals to retain in-house.
   3. 3.
     Framework Limitation: While CrypTen is designed to accommodate PyTorch, it is a research tool where not all plain text functionalities are implemented.
-Writing optimizers – such as L-BFGS – and custom operators that are not readily available requires both machine learning and cryptography knowledge.
+For example, writing optimizers – such as L-BFGS – and custom operators that are not readily available requires both machine learning and cryptography knowledge.
+Moreover, the protocol incurs additional computational overhead, especially if hyper-parameters become more complex to sweep [^4] For our work, the performance metrics are provided in Appendix 14 for reference.
+This will likely improve with time, as new frameworks address design shortcomings.
   4. 4.
     Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly.
-As detailed in Appendix  8, our hyperparameters for SGD are indeed different in encrypted and plaintext settings.
-However, encrypted computation hides loss curves and training details by default, complicating development.
+As detailed in Appendix 12, our hyperparameters for SGD differ in encrypted and plaintext settings.
+Thus, as encrypted computation hides loss curves and training details by default, development is expected to be complex.
+This is because both hospitals want to ensure model fit with secure evaluation, but may not want to expend the computational cost of private hyperparameter sweeping.
+Figure 9: Mean AUC ($\pm$ standard deviation).
+Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{s}(k=30)$ and $\pi_{0}$.
+Middle: All strategies perform similarly.
+Right: $\pi_{s}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
 8 Related Work
-5.1 Data Valuation and Pricing
-The question of how to assess the impact and therefore worth of data has been well studied.
-Data valuation as a field seeks to quantify the contribution of individual data points or datasets to model performance.
-Shapley value-based approaches provide theoretically grounded valuations but scale poorly to large datasets [45].
-More efficient methods include influence functions [46] and leave-one-out testing [47].
-Recent work has extended these concepts to dataset-level valuation [48].
-As practitioners create larger datasets, the emergence of data marketplaces has sparked interest in data pricing mechanisms [49].
-Query-based pricing [50] and outcome-driven valuations [51] aim to balance seller compensation with buyer utility.
-While these approaches inform fair data exchange, they typically assume direct access to data, unlike our privacy-preserving method.
-5.2 Alternative Approaches to Data Sharing
-Recent work has explored several approaches to mitigate data sharing constraints while maintaining model performance.
-We discuss two primary directions: synthetic data generation and transfer learning from public pretraining.
-Synthetic data generation
+We briefly relate alternative approaches towards dataset privacy for sharing. More work on data pricing, differential privacy, and private fine-tuning are included in Appendix 16.2.
+Augmenting existing data with synthetic data in medical domains
 Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy.
-Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data [52], medical imaging data [53], and electronic health records [54].
-These methods can preserve statistical properties of the original data while providing differential privacy guarantees.
-Transforming data into a similar form that de-sensitizes certain attributes can be desirable [55, 56, 57, 58, 59].
-Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself [60].
-Additionally, outside of the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks [61, 60].
-However, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables [52].
-For tabular data, methods like CTGAN [62] and TVAE [62] have demonstrated ability to learn complex distributions while preserving correlations between features.
-However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance [63].
-Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees [64].
-While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset [65].
-Public pre-training and private fine-tuning
-Transfer learning via public pretraining has become increasingly popular for domains with limited private data access.
-BioBERT [17] and ClinicalBERT [15] demonstrated that pretraining on PubMed abstracts and clinical notes can improve performance on downstream medical tasks.
-Similar approaches have emerged in other regulated domains, including FinBERT [66] for financial applications.
-However, the effectiveness of transfer learning depends heavily on domain alignment.
-One study showed that continued pretraining on domain-specific data significantly outperforms generic pretraining when domains differ substantially [16].
-This presents challenges for highly specialized fields where public data may not capture domain-specific patterns [67].
-Recent work has explored methods to quantify and optimize domain adaptation.
-Adaptive pretraining strategies [68] and domain-specific vocabulary augmentation [16] have shown promise in bridging domain gaps.
-However, these approaches still require substantial compute and may not fully capture specialized domain knowledge present in private datasets.
-5.3 Benefits of Data Scaling Beyond Performance
-Recent work has demonstrated that increasing dataset size and diversity yields benefits beyond raw performance metrics.
-Large-scale training data has been shown to improve model robustness to distribution shifts [23] and reduce demographic performance disparities [69].
-Studies of vision models trained on increasingly large datasets show improved out-of-distribution generalization [14].
-Similarly, language models trained on diverse data demonstrate better performance across different domains and demographic groups [70].
-5.4 Secure and Confidential Computation
-Secure and confidential computation encompasses cryptographic techniques that protect information privacy during computation.
-In a two-party setup between a model owner and data owner, these methods enable computing joint functions on private inputs without revealing them to other parties.
-This requires an encoding scheme $\mathrm{Enc}(\cdot)$ that satisfies the homomorphic property: $\mathrm{Enc}(A)\circ\mathrm{Enc}(B)=\mathrm{Enc}(A\circ B)$, where $A$ and $B$ represent data held by two parties.
-The inverse function $\mathrm{Enc}^{-1}(\cdot)$ must exist to decode the final output: $\mathrm{Enc}^{-1}(A\circ B)=A\circ B$.
-Considering an “honest-but-curious” threat model, where parties aim to jointly compute on privately-held data, two main approaches emerge.
-Fully Homomorphic Encryption (FHE)
-FHE enables arbitrary additions and multiplications on encrypted inputs.
-While it represents the gold standard for encrypted computation, adapting it to modern machine learning is challenging due to computational constraints from growing ciphertext size.
-FHE implementations typically use lattice-based schemes requiring periodic “bootstrapping” (key refreshing and noise reduction through re-encryption) via methods like the CKKS scheme [71].
-This introduces cryptographic parameters that non-experts struggle to configure effectively.
-Secure Multi-party Computation (S-MPC)
-SMPC enables multiple parties to compute functions over private inputs while revealing only the final output [33, 34].
-An MPC system uses key exchanges, encryption schemes, and secure communication to ensure only encrypted data leaves owner control [72].
-Though generally faster than FHE, SMPC’s engineering complexity and communication overhead can limit adoption.
-Traditional private training approaches that completely hide data can also impede essential model development tasks like inspection, monitoring, and debugging.
-5.5 Secure Data Combination
+Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data (Goncalves et al., 2020), medical imaging data (Thambawita et al., 2022), and electronic health records (Baowaly et al., 2019).
+These methods preserve statistical properties of the original data while providing differential privacy guarantees.
+Transforming data into a similar form that desensitizes certain attributes can be desirable (Drechsler, 2011; Howe et al., 2017; Nikolenko, 2021; Gonzales et al., 2023; Sweeney, 2002).
+Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself (Jordon et al., 2021).
+Additionally, outside the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks (Narayanan and Shmatikov, 2006; Jordon et al., 2021).
+In addition, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables (Goncalves et al., 2020).
+For tabular data, methods like CTGAN and TVAE (Xu et al., 2019) have demonstrated ability to learn complex distributions while preserving correlations between features.
+However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance (Assefa et al., 2020).
+Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees (Jordon et al., 2018).
+While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset (Yang et al., 2024a).
+Secure Data Combination
 Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance.
-Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data [73].
-However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets [74].
+Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data (Aono et al., 2017).
+However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets (McMahan et al., 2017).
 More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment.
-These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets [75, 76].
-Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information [77].
+These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets (Leung et al., 2019; Chakraborty et al., 2024).
+Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information (Duan et al., 2021).
+Others trained the downstream model in private, but limited to LASSO (van Egmond et al., 2021).
 Several systems have been developed to facilitate secure data combination in specific domains.
-In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance [78].
-Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality [66].
-5.6 Other Privacy-Preserving Methods
-Federated Learning.
-Cross-silo federated, decentralized, and collaborative machine learning  [74, 79, 80, 81] focus on acquiring more data through improved data governance and efficient system design.
-Healthcare machine learning is considered especially suitable, as health records are often isolated [82, 83, 84].
+In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance (Raisaro et al., 2018; van Egmond et al., 2021).
+Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality (Liu et al., 2021).
+Federated Learning
+Cross-silo federated, decentralized, and collaborative ML (McMahan et al., 2017; Li et al., 2020; Bonawitz et al., 2019; Kairouz et al., 2021) focus on acquiring more data through improved data governance and efficient system design.
+Healthcare machine learning is considered especially suitable, as health records are often isolated (Rieke et al., 2020; Xu et al., 2021; Nguyen et al., 2022; Cho et al., 2025).
 Yet, even though no raw data is shared, model parameters or gradients flow through the system.
-As the federated computing paradigm offer no privacy guarantee, the system is vulnerable to model inversion [85] and gradients leakage attacks [86, 87].
-A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimisation is traded off with privacy [88, 89].
-Building on the insight that useful data is often disparately owned, we tackle the specific incentive problem between pairs of data players where one side trains the model, instead of scaling up a federation (number of parties) to address data access issues.
-We thus focusing on making this exchange efficient, accurate, and private.
-Compared to vanilla Federated Learning, an MPC system [34, 33, 90, 35] provides stronger guarantee in terms of input security.
+As the federated computing paradigm offers no privacy guarantee, the system is vulnerable to model inversion (Geiping et al., 2020) and gradients leakage attacks (Boenisch et al., 2023; Zhu et al., 2019).
+A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimization is traded off with privacy (Lyu et al., 2022; Raynal and Troncoso, 2024).
+Building on the insight that useful data is often disparately-owned, we tackle the specific incentive problem between pairs of players where one side trains the model, instead of scaling up the number of parties through a federation.
+Compared to vanilla Federated Learning, an MPC system (Shamir, 1979; Yao, 1982; Bonawitz et al., 2017; Knott et al., 2021) provides stronger guarantee in terms of input security.
 Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements.
-Our work extends the line of works  [44, 91, 90] that demonstrates the potential of incorporating MPC in various federated scenarios.
-On the practical side, unlike mobile-based networks for secure federated learning protocols [80], our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
-Differential Privacy
-Differential privacy (DP) [92] offers formal privacy guarantees for sharing data and training machine learning models.
-While DP mechanisms can protect individual privacy when releasing model outputs or aggregated statistics, they face significant limitations for inter-organizational data sharing.
-The primary challenge is that DP operates on already-pooled data, but organizations are often unwilling to share their raw data in the first place [43].
-Even when organizations are willing to share data, the privacy guarantees of DP come at a substantial cost to utility, particularly in machine learning applications.
-DP-SGD, the standard approach for training deep neural networks with differential privacy, significantly degrades model performance compared to non-private training [93].
-This performance impact is especially pronounced in data-constrained settings, where recent work has shown that large models rely heavily on memorization of rare examples that DP mechanisms tend to obscure [94].
-The privacy-utility trade-off becomes even more challenging when dealing with high-dimensional data or complex learning tasks.
-Studies have demonstrated that achieving meaningful privacy guarantees while maintaining acceptable model performance requires prohibitively large datasets [95].
-This limitation is particularly problematic in specialized domains like healthcare, where data is inherently limited and performance requirements are stringent [96].
-Recent work has attempted to improve the privacy-utility trade-off through advanced composition theorems and adaptive privacy budget allocation [97].
-However, these approaches still struggle to match the performance of non-private training, especially when working with modern deep learning architectures [98].
-While differential privacy offers important theoretical guarantees, our work focuses on the practical challenge of enabling data owners to evaluate potential partnerships before sharing any data, addressing a key barrier to collaboration that DP alone cannot solve.
+Our work extends the line of work by  (Xu et al., 2022; Yang et al., 2024b; Bonawitz et al., 2017) that demonstrates the potential of incorporating MPC in various federated scenarios.
+On the practical side, unlike mobile-based networks for secure federated learning protocols (Bonawitz et al., 2019), our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
+Comparing with Federated Learning with Privacy Guarantees
+Privacy incentivizes federation (Usynin et al., 2024).
+Specifically, preserving privacy between the parties under federated learning uses secure computation methods (Truex et al., 2019; Bonawitz et al., 2017; Stripelis et al., 2021; Cho et al., 2025; Froelicher et al., 2021) combined with differential privacy (Xu et al., 2023; Usynin et al., 2024; Avent et al., 2017), trusted execution environments (Mo et al., 2021) – an approach that is coined ”Privacy-in-Depth” in Kairouz et al.
+(2021).
+Performing dataset evaluation with MPC, as in SecureKL, can be seen as an extension of data federation, by adding a separate privacy-preserving component before all parties commit to an entire federated learning system.
+It is low-commitment, as privacy is preserved by default, and there is no requirement to continue in any system; it tackles the incentives’ problem blocking collaboration.
+By demonstrating utility, the two parties foster trust.
+Our work complements the line of ambitious systems that incorporate MPC in potential federated scenarios where the participants are semi-honest (Yang et al., 2024b; Bonawitz et al., 2017; Cho et al., 2025).
 9 Conclusion
@@
 As the AI community continues to grapple with data access challenges, particularly in regulated domains like healthcare, methods that balance privacy and utility will become increasingly critical for responsible advancement of the field.
-As noted in Section 4.5, our approach has several limitations, including the fact that, despite impressive aggregate results, our method is less effective for individual hospitals; this finding is fertile ground for future work.
-Additionally, our work present opportunities for follow-up research.
+As noted in Section 6.3, our approach has several limitations, including the fact that, despite impressive aggregate results, our method is less effective for individual hospitals, which motivates future work.
 Our method assumes static datasets and may not generalize well to scenarios where data distributions evolve rapidly over time.